### v1

# The Fractional-Logarithmic Laplacian: Fundamental Properties and Eigenvalues

Abstract.
In this paper, we introduce, for the first time, the fractional–logarithmic Laplacian $(-\Delta)^{s+\operatorname*{Log}}$, defined as the derivative of the fractional Laplacian $(-\Delta)^{t}$ at $t=s$. It is a singular integral operator with Fourier symbol $|\xi|^{2s}(2\ln|\xi|)$ and we prove the pointwise integral representation
$(-\Delta)^{s+\operatorname*{Log}}=c_{n,s}\,\mathcal{\operatorname*{{\rm p.v.}}}\int_{\mathbb{R}^{n}}\frac{u(x)-u(y)}{|x-y|^{n+2s}}\bigl(-2\ln|x-y|\bigr)\,dy+b_{n,s}(-\Delta)^{s}u(x),$
where $c_{n,s}$ is the normalization constant of the fractional Laplacian, and $b_{n,s}:=\frac{d}{ds}c_{n,s}.$ We also establish several equivalent formulations of $(-\Delta)^{s+\operatorname*{Log}}$, including the singular-integral representation, the Fourier-multiplier representation, the spectral-calculus definition, and an extension characterization.
We develop the associated functional framework on both $\mathbb{R}^{n}$ and bounded Lipschitz domains, introducing the natural energy spaces and proving embedding results; in particular, we obtain a compact embedding at the critical exponent $2_{s}^{*}=\frac{2n}{n-2s}$, a phenomenon that differs from the classical Sobolev and fractional Sobolev settings. We further study the Poisson problem, proving existence and $L^{\infty}$-regularity results. We then investigate the Dirichlet eigenvalue problem and establish qualitative spectral properties. Finally, we derive a Weyl-type asymptotic law for the eigenvalue counting function and for the $k$-th Dirichlet eigenvalue, showing that the high-frequency behaviour combines the fractional Weyl scaling with the logarithmic growth factor, thus interpolating between the fractional Laplacian and the logarithmic Laplacian.
Key words and phrases:
Fractional-logarithmic Laplacian, Nonlocal operators, Poisson problem, Dirichlet eigenvalues, Weyl asymptotics
2020 Mathematics Subject Classification:
35B65, 47G20, 35A01, 35P20, 35R11
1. Introduction
In recent years, there has been a significant increase of interest in elliptic and parabolic boundary-value problems, as well as eigenvalue problems, involving hypersingular integral operators (see, for instance, [4, 19, 18, 22]). In this context, the fractional Laplacian $(-\Delta)^{s}$, for $0<s<1$, serves as a fundamental prototype and is defined by
(1.1) $(-\Delta)^{s}u(x)=c_{n,s}\,\operatorname*{{\rm p.v.}}\int_{\mathbb{R}^{n}}\frac{u(x)-u(y)}{|x-y|^{n+2s}}\,dy,$
for every $x\in\mathbb{R}^{n}$ and $u\in C_{c}^{2}(\mathbb{R}^{n})$. In (1.1), the normalization constant
(1.2) $c_{n,s}:=2^{2s}\pi^{-\frac{n}{2}}s\,\frac{\Gamma\!\left(\frac{n+2s}{2}\right)}{\Gamma(1-s)}$
is chosen so that $(-\Delta)^{s}$ has Fourier symbol $|\xi|^{2s}$ for $\xi\in\mathbb{R}^{n}$ (cf. [11, Chapter 3]), where $\Gamma$ denotes the Gamma function. With this normalization one also has the small-order limit
$\lim_{s\to 0^{+}}(-\Delta)^{s}u(x)=u(x)$
for $u\in C^{2}_{c}(\mathbb{R}^{n})$ and $x\in\mathbb{R}^{n}$.
Motivated by this limit, Weth and the first author [7] introduced the logarithmic Laplacian $\operatorname*{Log}(-\Delta)$ as the first-order correction term in the expansion as $s\to 0^{+}$. More precisely, for $u\in C^{\alpha}_{c}(\mathbb{R}^{n})$ with some $\alpha>0$, $\operatorname*{Log}(-\Delta)u$ is uniquely defined by
(1.3) $(-\Delta)^{s}u(x)=u(x)+s\,\operatorname*{Log}(-\Delta)u(x)+o(s)\ \ \text{in}\ \ L^{p}(\mathbb{R}^{n}),\ 1<p\leq\infty,\ s\rightarrow 0^{+}$
or equivalently,
$\operatorname*{Log}(-\Delta)u=\big.\frac{d}{ds}(-\Delta)^{s}u\big|_{s=0}\quad\text{in}\quad L^{p}(\mathbb{R}^{n}),\ 1<p\leq\infty.$
Moreover, they proved (see [7, Theorem 1.1]) that $\operatorname*{Log}(-\Delta)$ has the (weakly) singular Fourier symbol $2\ln|\xi|$ for $\xi\in\mathbb{R}^{n}$, and admits the integral representation
$\operatorname*{Log}(-\Delta)u(x)=c_{n}\operatorname*{{\rm p.v.}}\int_{B_{1}(0)}\frac{u(x)-u(x+y)}{|y|^{n}}dy-c_{n}\int_{\mathbb{R}^{n}\setminus B_{1}(0)}\frac{u(x+y)}{|y|^{n}}dy+\rho_{n}\,u(x),$
for every Dini-continuous function $u:\mathbb{R}^{n}\to\mathbb{R}$ belonging to the tail space
$L^{1}_{0}(\mathbb{R}^{n}):=\Bigl\{u\in L^{1}_{\mathrm{loc}}(\mathbb{R}^{n})\;:\;\int_{\mathbb{R}^{n}}(1+|x|)^{-n}|u(x)|\,dx<\infty\Bigr\}.$
Here $B_{1}(0)$ denotes the open unit ball in $\mathbb{R}^{n}$ centred at the origin and
$c_{n}=\frac{\Gamma(n/2)}{\pi^{n/2}}=\frac{2}{\omega_{n}},\qquad\rho_{n}:=2\ln 2+\psi(n/2)-\gamma,$
where $\gamma=-\Gamma^{\prime}(1)$ is the Euler–Mascheroni constant, $\psi=\Gamma^{\prime}/\Gamma$ is the digamma function, and $\omega_{n}:=|\mathbb{S}^{n-1}|$ denotes the (n1)(n-1)-dimensional surface measure of the unit sphere in $\mathbb{R}^{n}$.
More recently, the first and third authors, together with Weth [6], established that $\operatorname*{Log}(-\Delta)$ admits an extension property which is quite different from the Caffarelli-Silvestre extension for the fractional Laplacian (cf. [5]). Specifically, they showed that for every $u\in L^{1}_{0}(\mathbb{R}^{n})$,
$\operatorname*{Log}(-\Delta)u=2(\ln 2-\gamma)u-2\lim_{t\to 0^{+}}\Bigl(w_{u}(\cdot,t)+u\ln t\Bigr)\quad\text{in }\mathscr{D}^{\prime}(\mathbb{R}^{n}),$
where $w_{u}$ is the unique distributional solution to the Neumann problem
$\begin{cases}-\,{\rm div}\,_{(x,t)}(t\,\nabla w_{u})=0&\text{in }\mathbb{R}^{n+1}_{+},\\[2.0pt] -\displaystyle\lim_{t\to 0^{+}}t\,\partial_{t}w_{u}(\cdot,t)=0&\text{on }\mathbb{R}^{n},\end{cases}\qquad\mathbb{R}^{n+1}_{+}:=\mathbb{R}^{n}\times(0,\infty).$
Through various applications, it has become clear that it is useful to understand the logarithmic Laplacian $\operatorname*{Log}(-\Delta)$ and its fundamental properties. For instance, it appears in the study of the asymptotics of the Dirichlet (eigenvalue) problem for $(-\Delta)^{s}$ as $s\to 0^{+}$ (see [7, 13]); in $s$-dependent nonlinear Dirichlet problems in the small-order regime (see [1, 15]), motivated in particular by order-dependent optimization problems with small optimal order (see, e.g., [2, 21, 20] for applications in image processing and population dynamics); and in the geometric context of the 0-fractional perimeter (see [10]). Recently, the second author introduced in [8] a definition of $\operatorname*{Log}(-\Delta)$ on any complete Riemannian manifold using spectral-theoretic tools, together with an explicit integral representation. More recently, the conformal logarithmic Laplacian on the unit sphere has been studied by Fernández and Saldaña [12].
Returning to the expansion (1.3), it is natural to study the first-order derivative
$\big.\frac{d}{dt}(-\Delta)^{t}u\big|_{t=s}$
at an arbitrary order $0<s<1$ for some suitable function $u$. The following proposition establishes the existence of the derivative provided that $u\in C_{c}^{2}(\mathbb{R}^{n})$.
Proposition 1.1.
Let $0<s<1$ and $u\in C_{c}^{2}(\mathbb{R}^{n})$. Then for every fixed $x\in\mathbb{R}^{n}$, the map $t\mapsto(-\Delta)^{t}u(x)$ is $C^{1}$ on $(0,1)$, and
(1.4) $\left.\frac{d}{dt}(-\Delta)^{t}u(x)\right|_{t=s}$ $=c_{n,s}\,\mathcal{L}_{1}u(x)+b_{n,s}\,(-\Delta)^{s}u(x),$
where $c_{n,s}$ is given by (1.2),
(1.5) $b_{n,s}:=\frac{d}{ds}c_{n,s}=\ln 4+\frac{1}{s}+\psi(1-s)+\psi\!\left(\frac{n+2s}{2}\right).$
and
(1.6) $\mathcal{L}_{1}u(x):=\operatorname*{{\rm p.v.}}\int_{\mathbb{R}^{n}}\frac{u(x)-u(y)}{|x-y|^{n+2s}}\bigl(-2\ln|x-y|\bigr)\,dy,$
where $\psi$ denotes the digamma function, i.e., $\psi(z):=\frac{\Gamma^{\prime}(z)}{\Gamma(z)}$.
This motivates the following definition.
Definition 1.1.
For every $0<s<1$, we define the operator $(-\Delta)^{s+\operatorname*{Log}}:C_{c}^{2}(\mathbb{R}^{n})\to\mathbb{\mathbb{R}}$ by
$(-\Delta)^{s+\operatorname*{Log}}u(x):=\big.\frac{d}{dt}(-\Delta)^{t}u(x)\big|_{t=s},\quad x\in\mathbb{R}^{n}$
for every $u\in C_{c}^{2}(\mathbb{R}^{n})$ and call $(-\Delta)^{s+\operatorname*{Log}}$ the fractional-logarithmic Laplacian.
The aim of this paper is to study rigorously fundamental properties of the fractional-logarithmic Laplacian $(-\Delta)^{s+\operatorname*{Log}}$ on $\mathbb{R}^{n}$. Note, here and throughout this paper, we denote by $\mathcal{F}$ or $\widehat{\cdot}$ the Fourier transform
$\mathcal{F}[u](\xi)=\widehat{u}(\xi)=\frac{1}{(2\pi)^{n/2}}\int_{\mathbb{R}^{n}}e^{-i\xi\cdot x}\,u(x)\,dx,\qquad\xi\in\mathbb{R}^{n},$
for $u\in L^{1}(\mathbb{R}^{n})$. We also denote the Lizorkin space (see [24, Section 5.1.2])
$\mathcal{Z}(\mathbb{R}^{n}):=\Bigl\{\varphi\in\mathscr{S}(\mathbb{R}^{n}):\ \partial^{\alpha}\widehat{\varphi}(0)=0,\ \text{for all }\alpha\in\mathbb{N}_{0}^{\,n}\Bigr\},$
as well as the corresponding topological dual $\mathcal{Z}^{\prime}(\mathbb{R}^{n})$, i.e., the space of all continuous linear functionals on $\mathcal{Z}(\mathbb{R}^{n})$, where $\mathscr{S}(\mathbb{R}^{n})$ and $\mathscr{S}^{\prime}(\mathbb{R}^{n})$ denote, respectively, the Schwartz space on $\mathbb{R}^{n}$ and the space of tempered distributions on $\mathbb{R}^{n}$. With this choice, for every symbol $m(\xi)$ in
$m(\xi)\in\Bigl\{|\xi|^{2s},\ \ln|\xi|^{2},\ |\xi|^{2s}\ln|\xi|^{2}\Bigr\},$
the associated operator $m(D)$ defines a continuous linear map $\mathcal{Z}^{\prime}(\mathbb{R}^{n})\to\mathcal{Z}^{\prime}(\mathbb{R}^{n})$ by duality,
$\langle m(D)u,\varphi\rangle_{\mathcal{Z}^{\prime},\mathcal{Z}}:=\Big\langle u,\ \mathcal{F}^{-1}\!\big(m(\xi)\,\widehat{\varphi}(\xi)\big)\Big\rangle_{\mathcal{Z}^{\prime},\mathcal{Z}},\qquad u\in\mathcal{Z}^{\prime}(\mathbb{R}^{n}),\ \varphi\in\mathcal{Z}(\mathbb{R}^{n}).$
In particular, $(-\Delta)^{s}$, $\operatorname*{Log}(-\Delta)$ and $(-\Delta)^{s+\operatorname*{Log}}$ are well-defined continuous linear operators on $\mathcal{Z}^{\prime}(\mathbb{R}^{n})$.
Below we present several alternative representations and properties of the fractional-logarithmic Laplacian $(-\Delta)^{s+\operatorname*{Log}}$ ; depending on the setting, one may choose the most suitable formulation.
Theorem 1.1.
Let $s\in(0,1)$. Then the following statements hold.
  1. (i)
    (Integral representation) For every $u\in C_{c}^{2}(\mathbb{R}^{n})$ and $x\in\mathbb{R}^{n}$,
    $(-\Delta)^{s+\operatorname*{Log}}u(x)=c_{n,s}\,\mathcal{L}_{1}u(x)+b_{n,s}\,(-\Delta)^{s}u(x),$
    where $c_{n,s},b_{n,s},\mathcal{L}_{1}$ are defined in (1.2), (1.5) and (1.6), respectively.
  2. (ii)
    (Kernel representation) For every $u\in C_{c}^{2}(\mathbb{R}^{n})$ and $x\in\mathbb{R}^{n}$,
    $(-\Delta)^{s+\operatorname*{Log}}u(x)=\operatorname*{{\rm p.v.}}\int_{\mathbb{R}^{n}}\bigl(u(x)-u(y)\bigr)\,{\bf K}_{s+\operatorname*{Log}}(|x-y|)\,dy,$
    where
    (1.7) ${\bf K}_{s+\operatorname*{Log}}(r):=c_{n,s}\,\bigl(b_{n,s}-2\ln r\bigr)\,r^{-n-2s},\quad r>0.$
  3. (iii)
    (Fourier symbol) The fractional-logarithmic Laplacian $(-\Delta)^{s+\operatorname*{Log}}$ has the (weakly) singular Fourier symbol $|\xi|^{2s}\,(2\ln|\xi|)$ for $\xi\in\mathbb{R}^{n}$. More precisely,
    $\mathcal{F}\bigl[(-\Delta)^{s+\operatorname*{Log}}u\bigr](\xi)=|\xi|^{2s}\,\bigl(2\ln|\xi|\bigr)\,\widehat{u}(\xi),\qquad\xi\in\mathbb{R}^{n},$
    for every $u\in C_{c}^{2}(\mathbb{R}^{n})$.
  4. (iv)
    (Representation as a composition) For every $u\in\mathcal{Z}^{\prime}(\mathbb{R}^{n})$
    $(-\Delta)^{s+\operatorname*{Log}}u=(-\Delta)^{s}\bigl(\operatorname*{Log}(-\Delta)u\bigr)=\operatorname*{Log}(-\Delta)\bigl((-\Delta)^{s}u\bigr)\quad\text{in}\quad\mathcal{Z}^{\prime}(\mathbb{R}^{n}).$
  5. (v)
    (Stability) For every $u\in C_{c}^{2}(\mathbb{R}^{n})$,
    $(-\Delta)^{s+\operatorname*{Log}}u\to\operatorname*{Log}(-\Delta)u\quad\text{in }\quad L^{\infty}(\mathbb{R}^{n})\ \ \text{as }s\to 0^{+}.$
Concerning Theorem 1.1, we record the following remark.
Remark 1.1.
  1. (i)
    The limits of the constant $b_{n,s}$ in (1.5) satisfy
    $b_{n,s}\to+\infty\quad\text{as }s\to 0^{+},\qquad b_{n,s}\to-\infty\quad\text{as }s\to 1^{-}.$
    Moreover, $b_{n,s}$ is monotonically decreasing with respect to $s$, and there exists $s_{0}\in\left(\tfrac{1}{2},1\right)$ such that $b_{n,s}\geq 0$ for every $s\in(0,s_{0})$.
  2. (ii)
    The kernel ${\bf K}_{s+\operatorname*{Log}}$ in (1.7) is not nonnegative in general.
Below we present another perspective on the fractional-logarithmic Laplacian $(-\Delta)^{s+\operatorname*{Log}}$. In fact, functional calculus can provide an alternative and systematic framework for defining this operator. Starting from the spectral theorem, one defines spectral multipliers via the projection-valued spectral measure of the underlying self-adjoint operator. In the Euclidean setting this viewpoint becomes particularly transparent, since $\Delta$ is unitarily diagonalized by the Fourier transform and its spectral projections act as Fourier multipliers. As a consequence, the spectral-measure representation allows us to identify the functional-calculus definition of the corresponding spectral operators with the previously introduced definitions, thereby placing all of these constructions within a single unified theory, see [9].
By the spectral theorem, we can define the fractional Laplacian
$(-\Delta)^{s}=\int_{[0,\infty)}\lambda^{s}\,dE(\lambda):H^{2s}(\mathbb{R}^{n})\rightarrow L^{2}(\mathbb{R}^{n})$
with domain (spectral fractional space)
$H^{2s}(\mathbb{R}^{n}):=\Bigl\{u\in L^{2}(\mathbb{R}^{n}):\int_{[0,\infty)}\lambda^{2s}\,dE_{u,u}(\lambda)<\infty\Bigr\},$
see Section 2.2 for more details. Since
$\frac{\lambda^{t}-\lambda^{s}}{t-s}\;\xrightarrow{t\to s}\;\lambda^{s}\ln\lambda,\lambda>0,\ s\in(0,1),$
it is natural to define the fractional–logarithmic power of $\Delta$ by
$(-\Delta)_{\mathrm{spec}}^{s+\operatorname*{Log}}:=\int_{(0,\infty)}\lambda^{s}\ln\lambda\,dE(\lambda).$
More precisely,
$\bigl\langle(-\Delta)_{\mathrm{spec}}^{s+\operatorname*{Log}}u,v\bigr\rangle_{L^{2}}=\int_{(0,\infty)}\lambda^{s}\ln\lambda\,dE_{u,v}(\lambda),\quad u\in\operatorname{Dom}((-\Delta)_{\mathrm{spec}}^{s+\operatorname*{Log}}),\ v\in L^{2}(\mathbb{R}^{n}),$
where
(1.8) $\operatorname{Dom}((-\Delta)_{\mathrm{spec}}^{s+\operatorname*{Log}})=\Bigl\{u\in L^{2}(\mathbb{R}^{n}):\int_{(0,\infty)}\lambda^{2s}\ln^{2}\lambda\,dE_{u,u}(\lambda)<\infty\Bigr\}=:H^{2s+\operatorname*{Log}}(\mathbb{R}^{n}).$
It is easy to see that
$H^{s+\operatorname*{Log}}(\mathbb{R}^{n})\subset H^{s}(\mathbb{R}^{n})\subset H^{1}(\mathbb{R}^{n}),\ s\in(0,1).$
Proposition 1.2.
Fix $s\in(0,1)$ and assume that $u\in H^{2s_{0}}(\mathbb{R}^{n})$ for some $s_{0}>s.$ Then,
$\lim_{t\to s}\Bigl\|\frac{(-\Delta)^{t}-(-\Delta)^{s}}{t-s}\,u-(-\Delta)_{\mathrm{spec}}^{s+\operatorname*{Log}}u\Bigr\|_{L^{2}(\mathbb{R}^{n})}=0.$
Therefore, by Proposition 1.2 and the definition of fractional-logarithmic Laplacian, we can obtain that the spectral operator $(-\Delta)_{\mathrm{spec}}^{s+\operatorname*{Log}}$ coincides with $(-\Delta)^{s+\operatorname*{Log}}$ in $C_{c}^{2}(\mathbb{R}^{n})$. By the relation between the spectral measure of $\Delta$ and the Fourier transform (cf. [23, Chapter 7]):
$dE_{f,f}(\lambda)=F^{\prime}(\lambda)\,d\lambda=\frac{1}{2}\,\lambda^{\frac{n}{2}-1}\int_{\mathbb{S}^{n-1}}\bigl|\widehat{f}(\sqrt{\lambda}\,\omega)\bigr|^{2}\,d\omega\,d\lambda,$
and hence we obtain the following results:
Theorem 1.2.
Let $s>0$.
  1. (i)
    The spectral fractional–logarithmic Sobolev space defined in (1.8) admits the following characterization
    $H^{2s+\operatorname*{Log}}(\mathbb{R}^{n})=\Bigl\{u\in L^{2}(\mathbb{R}^{n}):\int_{\mathbb{R}^{n}}|\xi|^{4s}\,\ln^{2}\!\bigl(|\xi|^{2}\bigr)\,|\widehat{u}(\xi)|^{2}\,d\xi<\infty\Bigr\}.$
  2. (ii)
    For every $u\in H^{2s+\operatorname*{Log}}(\mathbb{R}^{n})$, $(-\Delta)^{s+\operatorname*{Log}}u$ is well-defined as an $L^{2}(\mathbb{R}^{n})$ function and satisfies
    (1.9) $(-\Delta)^{s+\operatorname*{Log}}u=\mathcal{F}^{-1}\!\Big(|\xi|^{2s}\ln(|\xi|^{2})\,\widehat{u}(\xi)\Big)\quad\text{in }\quad L^{2}(\mathbb{R}^{n}).$
    Moreover,
    $\|(-\Delta)^{s+\operatorname*{Log}}u\|_{L^{2}(\mathbb{R}^{n})}^{2}=\int_{\mathbb{R}^{n}}|\xi|^{4s}\,\ln^{2}\!\bigl(|\xi|^{2}\bigr)\,|\widehat{u}(\xi)|^{2}\,d\xi.$
Next we turn to the extension problem on the half-space $\mathbb{R}^{n+1}_{+}:=\mathbb{R}^{n}\times(0,\infty)$ of the fractional-logarithmic Laplacian $(-\Delta)^{s+\operatorname*{Log}}$. To do this, we introduce the following notation. For a given $u\in C_{c}^{\infty}(\mathbb{R}^{n})$, define
(1.10) $w_{s}(x,t)=p_{n,s}\int_{\mathbb{R}^{n}}\frac{t^{2s}u(y)}{(|x-y|^{2}+t^{2})^{\frac{n+2s}{2}}}\,dy,\qquad(x,t)\in\mathbb{R}^{n+1}_{+}.$
Then $w_{s}$ is the solution to the weighted Dirichlet problem
$\begin{cases}\,{\rm div}\,_{(x,t)}\!\big(t^{1-2s}\nabla_{(x,t)}w_{s}\big)=0&\text{in }\mathbb{R}^{n+1}_{+},\\ w_{s}(\cdot,0)=u&\text{on }\mathbb{R}^{n},\end{cases}$
where
$p_{n,s}:=\Big(\displaystyle\int_{\mathbb{R}^{n}}\frac{1}{(|z|^{2}+1)^{\frac{n+2s}{2}}}\,dz\Big)^{-1}=\pi^{-\frac{n}{2}}\frac{\Gamma(\frac{n+2s}{2})}{\Gamma(s)}.$
Set
(1.11) $v_{s}(x,t):=p_{n,s}\ \partial_{s}(\frac{w_{s}}{p_{n,s}})=p_{n,s}\ t^{2s}\int_{\mathbb{R}^{n}}\frac{2\ln t-\ln(|x-y|^{2}+t^{2})}{(|x-y|^{2}+t^{2})^{\frac{n+2s}{2}}}u(y)dy,$
and
(1.12) ${\bf d}_{s}:=\frac{c_{n,s}}{p_{n,s}}=2^{2s}\frac{s\ \Gamma(s)}{\Gamma(1-s)},\quad b_{1}:=p_{n,s}\int_{\mathbb{R}^{n}}\frac{-\ln(|z|^{2}+1)}{(|z|^{2}+1)^{\frac{n+2s}{2}}}dz.$
The following theorem provides an extension characterization of the fractional-logarithmic Laplacian.
Theorem 1.3.
Let $s\in(0,1),\ w_{s},v_{s},{\bf d}_{s},b_{1}$ be defined as in (1.10), (1.11) and (1.12) respectively and $u\in C^{\infty}_{c}(\mathbb{R}^{n})$. Then, $v_{s}$ is a solution of the inhomogeneous Dirichlet problem
(1.13) $\left\{\begin{array}[]{lll}\displaystyle\,{\rm div}\,_{(x,t)}\!\big(t^{1-2s}\nabla_{(x,t)}v_{s}\big)=2t^{-2s}\partial_{t}w_{s}\qquad&{\rm in}\ \ \mathbb{R}^{n+1}_{+},\\[7.11317pt] \phantom{----\!\!}\displaystyle v_{s}(\cdot,0)=b_{1}u\qquad&{\rm on}\ \ \mathbb{R}^{n}\end{array}\right$.
Moreover,
$(-\Delta)^{s+\operatorname*{Log}}u(x)$
$\quad=-\lim_{t\to 0^{+}}{\bf d}_{s}\Big[\big(b_{n,s}-2\ln t\big)t^{-2s}\big(w_{s}(x,t)-w_{s}(x,0)\big)+t^{-2s}\big(v_{s}(x,t)-v_{s}(x,0)\big)\Big].$
Here and in what follows, we denote $a_{\pm}:=\max\{0,\pm a\}.$ Set
${\bf k}_{s+\operatorname*{Log},\pm}(z):=c_{n,s}\,|z|^{-n-2s}\,(-\ln|z|)_{\pm},\qquad z\in\mathbb{R}^{n}\setminus\{0\},$
so that the full fractional–logarithmic kernel can be decomposed as
${\bf K}_{s+\operatorname*{Log}}(z)=c_{n,s}b_{n,s}\,|z|^{-n-2s}+2{\bf k}_{s+\operatorname*{Log},+}(z)-2{\bf k}_{s+\operatorname*{Log},-}(z).$
We denote by $\mathcal{H}^{s+\operatorname*{Log}}(\mathbb{R}^{n})$ the energy space associated with the positive part of the fractional–logarithmic kernel, namely
$\mathcal{H}^{s+\operatorname*{Log}}(\mathbb{R}^{n}):=\Bigl\{u\in L^{2}(\mathbb{R}^{n}):\ \iint_{\mathbb{R}^{n}\times\mathbb{R}^{n}}\bigl(u(x)-u(y)\bigr)^{2}\,{\bf k}_{s+\operatorname*{Log},+}(x-y)\,dx\,dy<\infty\Bigr\},$
where
${\bf k}_{s+\operatorname*{Log},+}(z):=c_{n,s}\,|z|^{-n-2s}\,(-\ln|z|)_{+},\qquad z\in\mathbb{R}^{n}\setminus\{0\}.$
For $u\in L^{2}(\mathbb{R}^{n})$, we define the Gagliardo-type seminorm
$u]_{s+\operatorname*{Log},+}^{2}:=\iint_{\mathbb{R}^{n}\times\mathbb{R}^{n}}\bigl(u(x)-u(y)\bigr)^{2}\,{\bf k}_{s+\operatorname*{Log},+}(x-y)\,dx\,dy.$
In particular,
$\mathcal{H}^{s+\operatorname*{Log}}(\mathbb{R}^{n})=\bigl\{u\in L^{2}(\mathbb{R}^{n}):\ [u]_{s+\operatorname*{Log},+}<\infty\bigr\},\ \|u\|_{\mathcal{H}^{s+\operatorname*{Log}}}^{\,2}:=\|u\|_{L^{2}(\mathbb{R}^{n})}^{\,2}+[u]_{s+\operatorname*{Log},+}^{2}.$
By Proposition 3.1, $\mathcal{H}^{s+\operatorname*{Log}}(\mathbb{R}^{n})$ is a Hilbert space. Next, we clarify the relationship between the spectral fractional–logarithmic Sobolev space and the energy space $\mathcal{H}^{s+\operatorname*{Log}}(\mathbb{R}^{n})$ defined here.
Proposition 1.3.
The embedding $H^{s+\operatorname*{Log}}(\mathbb{R}^{n})\hookrightarrow\mathcal{H}^{s+\operatorname*{Log}}(\mathbb{R}^{n})$ is continuous. In general the reverse inclusion fails, hence the embedding is typically strict. Moreover,
$\mathcal{H}^{s+\operatorname*{Log}}(\mathbb{R}^{n})=\Bigl\{u\in L^{2}(\mathbb{R}^{n}):\ \int_{\mathbb{R}^{n}}|\xi|^{2s}\ln(|\xi|^{2})\,|\widehat{u}(\xi)|^{2}\,d\xi<\infty\Bigr\}.$
In fact, this is consistent with the general principle for nonnegative self-adjoint operators: the operator domain is typically smaller than the form domain.
Let $\Omega\subset\mathbb{R}^{n}$ be a open set, we define the Dirichlet subspace by
$\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega):=\Bigl\{u\in\mathcal{H}^{s+\operatorname*{Log}}(\mathbb{R}^{n}):\ u=0\ \text{a.e. in }\Omega^{c}\Bigr\}.$
Throughout, we identify $L^{p}(\Omega)$ with the subspace of $L^{p}(\mathbb{R}^{n})$ consisting of functions vanishing a.e. on $\mathbb{R}^{n}\setminus\Omega$. Here and in what follows, we always assume that $\Omega\subset\mathbb{R}^{n}$ is a bounded Lipschitz domain.
We first establish a classical Poincaré inequality associated with our fractional–logarithmic operator.
Proposition 1.4.
If $u\in\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)$, then there exists a constant $C=C(n,s,\Omega)>0$ such that
(1.14) $u]_{s+\operatorname*{Log},+}^{2}\geq C\|u\|_{L^{2}(\Omega)}^{\,2}.$
As a consequence of Proposition 1.4, the seminorm $u]_{s+\operatorname*{Log},+}$ defines a norm on $\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)$. Define
$\|u\|_{\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)}:=[u]_{s+\operatorname*{Log},+}.$
Moreover, $\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)$ is a closed subspace of $\mathcal{H}^{s+\operatorname*{Log}}(\mathbb{R}^{n})$ and therefore a Hilbert space. For $u,w\in C_{c}^{\infty}(\Omega)$, we introduce the bilinear form
$\mathcal{E}_{\pm}(u,w):=\iint_{\mathbb{R}^{n}\times\mathbb{R}^{n}}(u(x)-u(y))(w(x)-w(y))\,{\bf k}_{s+\operatorname*{Log},\pm}(x-y)\,dx\,dy,$
and
(1.15) $\mathcal{E}_{s+\operatorname*{Log}}(u,w):=\mathcal{E}_{+}(u,w)-\mathcal{E}_{-}(u,w)+\frac{b_{n,s}c_{n,s}}{2}\,\mathcal{E}_{s}(u,w),$
where
$\mathcal{E}_{s}(u,w):=\iint_{\mathbb{R}^{n}\times\mathbb{R}^{n}}(u(x)-u(y))(w(x)-w(y))\,|x-y|^{-n-2s}\,dx\,dy,$
so that
$\|u\|_{\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)}=[u]_{s+\operatorname*{Log},+}=\sqrt{\mathcal{E}_{+}(u,u)}.$
The full energy $\mathcal{E}_{s+\operatorname*{Log}}$ is not always positive definite. The following proposition provides a more refined analysis of the bilinear forms introduced above.
Proposition 1.5.
Let $s\in(0,1)$ and $\Omega$ be bounded.
  1. (i)
    The negative part satisfies the estimate
    $0\leq\mathcal{E}_{-}(u,u)\leq\frac{1}{s^{2}}\,c_{n,s}\,|\mathbb{S}^{n-1}|\,\|u\|_{L^{2}(\Omega)}^{2},\quad u\in L^{2}(\Omega),$
    where $c_{n,s}$ is given by (1.2).
  2. (ii)
    Let $u\in\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)$ and set
    $\operatorname{diam}(\Omega):=\sup\{|x-y|:\ x,y\in\Omega\}<1.$
    Then
    (1.16) $\mathcal{E}_{+}(u,u)-\mathcal{E}_{-}(u,u)\geq\frac{1}{s}\,c_{n,s}\,|\mathbb{S}^{n-1}|\,\operatorname{diam}(\Omega)^{-2s}\Bigl(\ln\frac{1}{\operatorname{diam}(\Omega)}-\frac{1}{2s}\Bigr)\,\|u\|_{L^{2}(\Omega)}^{\,2}.$
    In particular, if $\operatorname{diam}(\Omega)\leq e^{-1/(2s)}$, then
    $\mathcal{E}_{+}(u,u)-\mathcal{E}_{-}(u,u)\geq 0.$
  3. (iii)
    Assume that $\operatorname{diam}(\Omega)<e^{-1/(2s)}$ and $b_{n,s}\geq 0$. Then $\mathcal{E}_{s+\operatorname*{Log}}(u,u)$ is positive definite on $\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)$. Moreover, $|u|\in\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)$ and
    (1.17) $\mathcal{E}_{s+\operatorname*{Log}}(|u|,|u|)\leq\mathcal{E}_{s+\operatorname*{Log}}(u,u).$
    Equality holds in (1.17) if and only if $u$ does not change sign.
Therefore, the quadratic form $\mathcal{E}_{s+\operatorname*{Log}}$ associated with $(-\Delta)^{s+\operatorname*{Log}}$ is well-defined on $\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)$. We recall that $\mathcal{H}^{s}_{0}(\Omega)$ is the space of measurable functions $u:\mathbb{R}^{n}\to\mathbb{R}$ such that $u\equiv 0$ on $\mathbb{R}^{n}\setminus\Omega$ and
$\|u\|_{\mathcal{H}^{s}_{0}(\Omega)}:=\Biggl(\iint_{\mathbb{R}^{n}\times\mathbb{R}^{n}}\frac{(u(x)-u(y))^{2}}{|x-y|^{n+2s}}\,dx\,dy\Biggr)^{1/2}<\infty.$
Equipped with this norm, $\mathcal{H}^{s}_{0}(\Omega)$ is a Hilbert space. It is well known that (see [18, Part I])
  1. (i)
    (Continuous embedding) If $n>2s$ and $2_{s}^{*}:=\frac{2n}{n-2s}$, then
    $\|u\|_{L^{2_{s}^{*}}(\Omega)}\leq C\,\|u\|_{\mathcal{H}^{s}_{0}(\Omega)}\qquad\text{for all }u\in\mathcal{H}^{s}_{0}(\Omega).$
    More generally, for every $q\in[1,2_{s}^{*}]$ one has a continuous embedding $\mathcal{H}^{s}_{0}(\Omega)\hookrightarrow L^{q}(\Omega)$:
    $\|u\|_{L^{q}(\Omega)}\leq C_{q}\,\|u\|_{\mathcal{H}^{s}_{0}(\Omega)}.$
  2. (ii)
    (Compact embedding) If $n>2s$, then the embedding $\mathcal{H}^{s}_{0}(\Omega)\hookrightarrow L^{q}(\Omega)$ is compact for every
    $1\leq q<2_{s}^{*}.$
Next, we clarify the relation between the fractional–logarithmic space and the standard fractional Sobolev spaces. In particular, $\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)$ is an intermediate space between $\mathcal{H}^{s+\varepsilon}_{0}(\Omega)$ and $\mathcal{H}^{s}_{0}(\Omega)$ for every $\varepsilon\in(0,1-s).$ Moreover, at the critical exponent $2_{s}^{*}$, we still obtain compactness of the embedding into $L^{2_{s}^{*}}(\Omega)$. This is markedly different from the classical Sobolev and fractional Sobolev settings. The key mechanism is the stronger singular behavior induced by the logarithmic factor, whose contribution is essential in the proof. We provide a concise compactness argument, with Lemma 3.1 as the main tool.
Proposition 1.6.
Let $s\in(0,1)$ and $n>2s.$
  1. (i)
    For $\varepsilon\in(0,1-s)$, one has the continuous inclusion
    $\mathcal{H}^{s+\varepsilon}_{0}(\Omega)\subset\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)\subset\mathcal{H}^{s}_{0}(\Omega).$
  2. (ii)
    The embedding $\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)\hookrightarrow L^{p}(\Omega)$ is compact for every $p\in[1,2_{s}^{*}]$.
In particular, we also study the existence, uniqueness, and regularity properties of weak solutions $u$ to the nonlocal Poisson problem
(1.18) $(-\Delta)^{s+\operatorname*{Log}}u+V(x)\,u$ $=f\qquad$ $\text{in }\Omega,$
$u$ $=0\qquad$ $\text{in }\Omega^{c},$
driven by the fractional–logarithmic operator $(-\Delta)^{s+\operatorname*{Log}}$, where $V\in L^{\infty}(\Omega)$.
Fix $r\in(0,e^{-|b_{n,s}|/2})$ and set
(1.19) $\alpha_{r}:=1-\frac{|b_{n,s}|}{2\,\ln(1/r)}>0.$
Denote $\operatorname{essinf}_{\Omega}V$ the essential infimum of $V$ on $\Omega$, i.e.
$\operatorname{essinf}_{\Omega}V=\sup\{a\in\mathbb{R}:\ V(x)\geq a\ \text{for a.e. }x\in\Omega\}.$
We now apply the Lax–Milgram theorem to establish existence (and uniqueness) of weak solutions to the Poisson problem, and then use a Moser iteration scheme to derive $L^{\infty}$-regularity.
Theorem 1.4.
Let $\Omega\subset\mathbb{R}^{n}$ be bounded, and consider the Poisson problem (1.18).
  1. (i)
    Assume that
    (1.20) $\operatorname*{ess\,inf}_{\Omega}V\ \geq\ \frac{c_{n,s}|\mathbb{S}^{n-1}|}{s^{2}}+\frac{|b_{n,s}|\,c_{n,s}}{s}\,|\mathbb{S}^{n-1}|\,r^{-2s}.$
    Then (1.18) admits a unique weak solution $u\in\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega).$ Moreover, the solution satisfies
    $\|u\|_{\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)}\leq\frac{1}{\alpha_{r}}\,\|f\|_{(\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega))^{*}},$
    where $c_{n,s},b_{n,s},\alpha_{r}$ are defined in (1.2), (1.5) and (1.19) respectively.
  2. (ii)
    Let $u$ be a weak solution of (1.18). If $f\in L^{q}(\Omega)$ for some $q>\frac{n}{2s},$ then $u\in L^{\infty}(\Omega).$
Our next aim is to study the associated Dirichlet eigenvalue problem
(1.21) $\left\{\begin{aligned} (-\Delta)^{s+\operatorname*{Log}}u&=\lambda\,u\qquad&&\text{in }\Omega,\\ u&=0\qquad&&\text{in }\mathbb{R}^{n}\setminus\Omega,\end{aligned}\right$.
where $\Omega$ is a bounded Lipschitz domain in $\mathbb{R}^{n}$.
We say that $u\in\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)$ is an eigenfunction of (1.21) associated with the eigenvalue $\lambda$ if
$\mathcal{E}_{s+\operatorname*{Log}}(u,\varphi)=\lambda\int_{\Omega}u\,\varphi\,dx\qquad\text{for all }\varphi\in\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega).$
Theorem 1.5.
Let $\Omega$ be a bounded Lipschitz domain in $\mathbb{R}^{n}$. Then problem (1.21) admits a nondecreasing sequence of real eigenvalues
$\lambda_{1}^{s+\operatorname*{Log}}(\Omega)\leq\lambda_{2}^{s+\operatorname*{Log}}(\Omega)\leq\cdots\leq\lambda_{k}^{s+\operatorname*{Log}}(\Omega)\leq\cdots,$
with corresponding eigenfunctions $\{\xi_{k}\}_{k\in\mathbb{N}}\subset\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)$ such that:
  1. (i)
    For every $k\in\mathbb{N}$,
    $\lambda_{k}^{s+\operatorname*{Log}}(\Omega)=\min\Bigl\{\mathcal{E}_{s+\operatorname*{Log}}(u,u)\ :\ u\in\mathcal{H}_{k}(\Omega),\ \|u\|_{L^{2}(\Omega)}=1\Bigr\},$
    where $\mathcal{H}_{1}(\Omega):=\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)$ and, for $k\geq 2$,
    $\mathcal{H}_{k}(\Omega):=\Bigl\{u\in\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)\ :\ \int_{\Omega}u\,\xi_{i}\,dx=0\ \text{for }i=1,\dots,k-1\Bigr\}.$
  2. (ii)
    One has $\lambda_{k}^{s+\operatorname*{Log}}(\Omega)\to+\infty$ as $k\to\infty$.
  3. (iii)
    The family $\{\xi_{k}\}_{k\in\mathbb{N}}$ forms an orthonormal basis of $L^{2}(\Omega)$.
  4. (iv)
    Assume that $\operatorname{diam}(\Omega)<e^{-1/(2s)}$ and $b_{n,s}\geq 0$. Then
    $\lambda_{1}^{s+\operatorname*{Log}}(\Omega)>0,$
    and the first eigenfunction $\xi_{1}$ can be chosen nonnegative.
Finally, we derive the Weyl-type asymptotic law for the eigenvalues of the fractional–logarithmic operator.
Theorem 1.6.
Let $\Omega$ is a bounded Lipschitz domain in $\mathbb{R}^{n}$, and let $\{\lambda_{k}^{s+\operatorname*{Log}}(\Omega)\}_{k\in\mathbb{N}}$ be the eigenvalues of (1.21). For $\Lambda>0$ define the counting function
$\mathcal{N}(\Lambda):=\#\bigl\{k\in\mathbb{N}:\ \lambda_{k}^{s+\operatorname*{Log}}(\Omega)\leq\Lambda\bigr\}=\sum_{k\in\mathbb{N}}(\Lambda-\lambda_{k})_{+}^{0}.$
Then
$\lim_{\Lambda\to+\infty}\,\mathcal{N}(\Lambda)\,\Lambda^{-\frac{n}{2s}}\bigl(\ln\Lambda\bigr)^{\frac{n}{2s}}=(2\pi)^{-n}\,s^{\frac{n}{2s}}\,\omega_{n}\,|\Omega|,$
and
$\lim_{k\to+\infty}\frac{\lambda_{k}^{s+\operatorname*{Log}}(\Omega)\,k^{-\frac{2s}{n}}}{\ln k}=\frac{2}{n}\,(2\pi)^{2s}\bigl(\omega_{n}|\Omega|\bigr)^{-\frac{2s}{n}},$
where $\omega_{n}$ denotes the volume of the unit ball in $\mathbb{R}^{n}$.
Remark 1.2.
For the fractional Laplacian $(-\Delta)^{s}$ on a bounded Lipschitz domain $\Omega\subset\mathbb{R}^{n}$, the eigenvalues $\{\lambda_{k}^{s}(\Omega)\}_{k\in\mathbb{N}}$ satisfy the Weyl asymptotics
(1.22) $\lim_{k\to\infty}\lambda_{k}^{s}(\Omega)\,k^{-\frac{2s}{n}}=(2\pi)^{2s}\bigl(\omega_{n}|\Omega|\bigr)^{-\frac{2s}{n}}.$
For the logarithmic Laplacian $\operatorname*{Log}(-\Delta)$, the eigenvalues $\{\lambda^{\ln}_{k}(\Omega)\}_{k\in\mathbb{N}}$ satisfy
(1.23) $\lim_{k\to\infty}\frac{\lambda^{\ln}_{k}(\Omega)}{\ln k}=\frac{2}{n}.$
Consequently, the operator $(-\Delta)^{s+\operatorname*{Log}}$ interpolates between the fractional and logarithmic regimes, and the asymptotic law in Theorem 1.6 is precisely the product of the fractional Weyl constant in (1.22) and the logarithmic growth factor in (1.23).
The remainder of this paper is organized as follows. In Section 2, we present and justify several equivalent formulations of the fractional-logarithmic operator, including the singular-integral representation, the Fourier-multiplier representation, the spectral-calculus definition, and the extension characterization. In Section 3, we first establish key properties of the underlying function spaces and energy forms, and then prove existence and regularity results for the Poisson problem. In Section 4, we study the Dirichlet eigenvalue problem, including qualitative properties of eigenvalues and the Weyl-type asymptotic law.
2. Representation Formulas and Basic Properties
In this section, we provide proofs of the various characterizations of the fractional-logarithmic operator, including the singular-integral representation, the Fourier-multiplier representation, the spectral-calculus definition, and the extension characterization.
2.1. Equivalent Represnetations
We begin with the proof of Proposition 1.1, which establishes the existence of the derivative of the fractional Laplacian with respect to the order at a general $s\in(0,1)$.
Proof of Proposition 1.1..
Since $u\in C_{c}^{2}(\mathbb{R}^{n})$, we have the Taylor expansion
$u(y)=u(x)+\nabla u(x)\cdot(y-x)+O(|x-y|^{2})\quad\text{as}\qquad y\to x$
Hence
$u(x)-u(y)=-\nabla u(x)\cdot(y-x)+O(|x-y|^{2}).$
In the principal value sense, the contribution of the linear term vanishes by symmetry: for every $\varepsilon>0$,
$\int_{|y-x|>\varepsilon}\frac{\nabla u(x)\cdot(x-y)}{|x-y|^{n+2t}}\,dy=0.$
Therefore, near $y=x$ the integrand behaves like
$\frac{O(|x-y|^{2})}{|x-y|^{n+2t}},$
which is locally integrable for all $t\in(0,1)$. This shows that $(-\Delta)^{t}u(x)$ is well-defined for every $t\in(0,1)$.
Fix $x\in\mathbb{R}^{n}$ and write
$I(t):=\operatorname*{{\rm p.v.}}\int_{\mathbb{R}^{n}}\frac{u(x)-u(y)}{|x-y|^{n+2t}}\,dy,\qquad\text{so that}\qquad(-\Delta)^{t}u(x)=c_{n,t}I(t).$
Moreover,
$\partial_{t}\Big(|x-y|^{-n-2t}\Big)=-2\ln|x-y|\,|x-y|^{-n-2t}.$
Splitting the integral into $B_{1}(x)$ and $\mathbb{R}^{n}\setminus B_{1}(x)$, the far-field part is absolutely integrable for all $t\in(0,1)$ since $u$ has compact support. For the near-field part, using the Taylor expansion $u(x)-u(y)=-\nabla u(x)\cdot(y-x)+O(|x-y|^{2})$ as $y\to x$ and the symmetry of the principal value, the linear term cancels and the remaining integrand behaves like $|x-y|^{-n-2t+2}|\ln|x-y||$, which is integrable near $y=x$ for every $t\in(0,1)$. Hence $I$ is $C^{1}$ and
$I^{\prime}(s)=\operatorname*{{\rm p.v.}}\int_{\mathbb{R}^{n}}\frac{u(x)-u(y)}{|x-y|^{n+2s}}\bigl(-2\ln|x-y|\bigr)\,dy.$
Finally,
$\left.\frac{d}{dt}(-\Delta)^{t}u(x)\right|_{t=s}$ $=c_{n,s}^{\prime}I(s)+c_{n,s}I^{\prime}(s)$
$=\frac{c_{n,s}^{\prime}}{c_{n,s}}(-\Delta)^{s}u(x)-2c_{n,s}\operatorname*{{\rm p.v.}}\int_{\mathbb{R}^{n}}\frac{(u(x)-u(y))\ln|x-y|}{|x-y|^{n+2s}}\,dy,$
which is (1.4). Recall that
$c_{n,s}=2^{2s}\pi^{-\frac{n}{2}}s\,\frac{\Gamma\!\left(\frac{n+2s}{2}\right)}{\Gamma(1-s)}.$
Taking logarithms gives
$\ln c_{n,s}=2s\ln 2-\frac{n}{2}\ln\pi+\ln s+\ln\Gamma\!\left(\frac{n+2s}{2}\right)-\ln\Gamma(1-s).$
Differentiating with respect to $s$ and using $\psi=\Gamma^{\prime}/\Gamma$ (the digamma function), we obtain
$b_{n,s}=\frac{c_{n,s}^{\prime}}{c_{n,s}}=\frac{d}{ds}\ln c_{n,s}=2\ln 2+\frac{1}{s}+\psi\!\left(\frac{n+2s}{2}\right)+\psi(1-s).$
We now present proofs of further representations of the fractional-logarithmic operator.
Proof of Theorem 1.1..
Parts (i) and (ii) follow directly from Proposition 1.1.
(iii) It follows by [11, Chapter 3] that
$c_{n,s}=\Big(\int_{\mathbb{R}^{n}}\frac{1-\cos\zeta_{1}}{|\zeta|^{n+2s}}d\zeta\Big)^{-1}.$
For $s\in(0,1)$, we obtain that
$\Big|\frac{u(x+y)+u(x-y)-2u(x)}{|y|^{n+2s}}\big(-2\log|y|\big)\Big|$
$\qquad\leq 2\chi_{B_{1}(y)}|y|^{-n-2s+2}\big|\log|y|\big|\|u\|_{C^{2}(\mathbb{R}^{n})}$
$\qquad\qquad+2\chi_{\mathbb{R}^{n}\setminus B_{1}(y)}|y|^{-n-2s}\big|\log|y|\big|\big|u(x+y)+u(x-y)-2u(x)\big|,$
then $\frac{u(x+y)+u(x-y)-2u(x)}{|y|^{n+2s}}\big(-2\log|y|\big)\in L^{1}(\mathbb{R}^{2n})$. Consequently, by the Fubini-Tonelli theorem, we can exchange the integral in $y$ with the Fourier transform in $x$. Thus, we apply the Fourier transform in the variable $x$ and we obtain
${\mathcal{F}}\big(c_{n,s}{\mathcal{L}}_{1}u\big)(\xi)$ $=-\frac{c_{n,s}}{2}\int_{\mathbb{R}^{n}}\frac{e^{i\xi\cdot y}+e^{-i\xi\cdot y}-2}{|y|^{n+2s}}\big(\!-2\log|y|\big)dy\,\hat{u}(\xi)$
$=c_{n,s}\int_{\mathbb{R}^{n}}\frac{1-\cos(\xi\cdot y)}{|y|^{n+2s}}\big(\!-2\log|y|\big)dy\,\hat{u}(\xi),$
where
$\int_{\mathbb{R}^{n}}\frac{1-\cos(\xi\cdot y)}{|y|^{n+2s}}\big(\!\!-2\log|y|\big)dy$
$=$ $\int_{\mathbb{R}^{n}}\frac{1-\cos(|\xi|y_{1})}{|y|^{n+2s}}\big(\!-2\log|y|\big)dy$
$=$ $|\xi|^{2s}\int_{\mathbb{R}^{n}}\frac{1-\cos z_{1}}{|z|^{n+2s}}\big(2\log|\xi|-2\log|z|\big)dz$
$=$ $|\xi|^{2s}\big(2\log|\xi|\big)\int_{\mathbb{R}^{n}}\frac{1-\cos z_{1}}{|z|^{n+2s}}dz+|\xi|^{2s}\int_{\mathbb{R}^{n}}\frac{1-\cos z_{1}}{|z|^{n+2s}}\big(\!-2\log|z|\big)dz$
$=$ $c_{n,s}^{-1}|\xi|^{2s}\big(2\log|\xi|\big)+|\xi|^{2s}\int_{\mathbb{R}^{n}}\frac{1-\cos z_{1}}{|z|^{n+2s}}\big(\!-2\log|z|\big)dz.$
Thus, we obtain that
${\mathcal{F}}((-\Delta)^{s+\operatorname*{Log}}u)(\xi)={\mathcal{F}}\big(c_{n,s}{\mathcal{L}}_{1}u+\widehat{b}_{n,s}(-\Delta)^{s}u\big)(\xi)=|\xi|^{2s}\big(2\log|\xi|\big)\hat{u}(\xi),$
where
$\widehat{b}_{n,s}=c_{n,s}\int_{\mathbb{R}^{n}}\frac{1-\cos\zeta_{1}}{|\zeta|^{n+2s}}\big(2\log(|\zeta|)\big)d\zeta.$
It is sufficient to prove that $\widehat{b}_{n,s}=b_{n,s}.$
Split $\mathbb{R}^{n}=B_{1}(0)\cup(\mathbb{R}^{n}\setminus B_{1}(0))$. As $|\zeta|\to 0$,
$1-\cos\zeta_{1}\sim\frac{\zeta_{1}^{2}}{2}\lesssim|\zeta|^{2},$
hence
$\frac{1-\cos\zeta_{1}}{|\zeta|^{n+2s}}\;|\ln|\zeta||\;\lesssim\;|\zeta|^{-n-2s+2}\,|\ln|\zeta||,$
which is integrable on $B_{1}(0)$ for $s\in(0,1)$ since in polar coordinates it becomes $\int_{0}^{1}r^{1-2s}|\ln r|\,dr<\infty$. As $|\zeta|\to\infty$, we have $0\leq 1-\cos\zeta_{1}\leq 2$, so
$\frac{1-\cos\zeta_{1}}{|\zeta|^{n+2s}}\,|\ln|\zeta||\;\lesssim\;|\zeta|^{-n-2s}\ln|\zeta|,$
which is integrable on $\mathbb{R}^{n}\setminus B_{1}(0)$ for any $s>0$. Therefore, by dominated convergence we may differentiate under the integral sign and obtain
$\frac{d}{ds}c_{n,s}^{-1}=\int_{\mathbb{R}^{n}}(1-\cos\zeta_{1})\,\frac{d}{ds}\big(|\zeta|^{-n-2s}\big)\,d\zeta=\int_{\mathbb{R}^{n}}\frac{1-\cos\zeta_{1}}{|\zeta|^{n+2s}}\bigl(-2\ln|\zeta|\bigr)\,d\zeta=\widehat{b}_{n,s}.$
Hence $\widehat{b}_{n,s}=\frac{c_{n,s}^{\prime}}{c_{n,s}}=b_{n,s}$.
(iv) Recall that, for a symbol $m(\xi)$, the corresponding operator $m(D)$ on $\mathcal{Z}^{\prime}(\mathbb{R}^{n})$ is defined by duality via
(2.1) $\langle m(D)u,\varphi\rangle_{\mathcal{Z}^{\prime},\mathcal{Z}}:=\Big\langle u,\ \mathcal{F}^{-1}\!\big(m(\xi)\widehat{\varphi}(\xi)\big)\Big\rangle_{\mathcal{Z}^{\prime},\mathcal{Z}},\qquad u\in\mathcal{Z}^{\prime}(\mathbb{R}^{n}),\ \varphi\in\mathcal{Z}(\mathbb{R}^{n}).$
In particular,
$(-\Delta)^{s}\equiv m_{s}(D)\ \text{with}\ m_{s}(\xi)=|\xi|^{2s},\qquad\operatorname*{Log}(-\Delta)\equiv\ell(D)\ \text{with}\ \ell(\xi)=\operatorname*{Log}|\xi|^{2},$
and
$(-\Delta)^{s+\operatorname*{Log}}\equiv m_{s+\operatorname*{Log}}(D)\ \text{with}\ m_{s+\operatorname*{Log}}(\xi)=|\xi|^{2s}\operatorname*{Log}|\xi|^{2}.$
Fix $u\in\mathcal{Z}^{\prime}(\mathbb{R}^{n})$ and $\varphi\in\mathcal{Z}(\mathbb{R}^{n})$. Using (2.1) twice and the fact that multiplication of symbols is commutative, we compute
$\big\langle(-\Delta)^{s}(\operatorname*{Log}(-\Delta)u),\varphi\big\rangle_{\mathcal{Z}^{\prime},\mathcal{Z}}$ $=\big\langle\operatorname*{Log}(-\Delta)u,\ \mathcal{F}^{-1}\!\big(|\xi|^{2s}\widehat{\varphi}(\xi)\big)\big\rangle_{\mathcal{Z}^{\prime},\mathcal{Z}}$
$=\Big\langle u,\ \mathcal{F}^{-1}\!\Big(\operatorname*{Log}|\xi|^{2}\cdot|\xi|^{2s}\widehat{\varphi}(\xi)\Big)\Big\rangle_{\mathcal{Z}^{\prime},\mathcal{Z}}$
$=\Big\langle u,\ \mathcal{F}^{-1}\!\Big(|\xi|^{2s}\operatorname*{Log}|\xi|^{2}\,\widehat{\varphi}(\xi)\Big)\Big\rangle_{\mathcal{Z}^{\prime},\mathcal{Z}}$
$=\big\langle(-\Delta)^{s+\operatorname*{Log}}u,\varphi\big\rangle_{\mathcal{Z}^{\prime},\mathcal{Z}}.$
An analogous computation yields
$\big\langle\operatorname*{Log}(-\Delta)((-\Delta)^{s}u),\varphi\big\rangle_{\mathcal{Z}^{\prime},\mathcal{Z}}=\Big\langle u,\ \mathcal{F}^{-1}\!\Big(\operatorname*{Log}|\xi|^{2}\cdot|\xi|^{2s}\widehat{\varphi}(\xi)\Big)\Big\rangle_{\mathcal{Z}^{\prime},\mathcal{Z}}=\big\langle(-\Delta)^{s+\operatorname*{Log}}u,\varphi\big\rangle_{\mathcal{Z}^{\prime},\mathcal{Z}}.$
Since $\varphi\in\mathcal{Z}(\mathbb{R}^{n})$ is arbitrary, the identities hold in $\mathcal{Z}^{\prime}(\mathbb{R}^{n})$.
(v) Since $u\in C^{2}_{c}(\mathbb{R}^{n})$ and $s\in(0,\frac{1}{4})$, then for some $C_{0}>0$
$|u(x)-u(y)|\leq C_{0}\min\{|x-y|,1\},$
and some $R>2$, supp $u\subset B_{R}(0)$. Let
$d_{n,s}=2^{2s}\pi^{-\frac{n}{2}}\frac{\Gamma(\frac{n+2s}{2})}{\Gamma(1-s)},$
which is bounded and
$d_{n,s}\to\pi^{-\frac{n}{2}}\Gamma(\frac{n}{2})\quad{\rm as}\ \,s\to 0^{+}.$
Moreover
$c_{n,s}=2^{2s}\pi^{-\frac{n}{2}}s\frac{\Gamma(\frac{n+2s}{2})}{\Gamma(1-s)}=sd_{n,s}$
and
$c_{n,s}b_{n,s}=d_{n,s}\Big[1+\Big(\ln 4+\psi(1-s)+\psi(\frac{n+2s}{2})\Big)s\Big]\to\pi^{-\frac{n}{2}}\Gamma(\frac{n}{2})=c_{n}\quad{\rm as}\ \,s\to 0^{+}.$
Part 1: estimates for $c_{n,s}{\mathcal{L}}_{1}u(x)$: Observe that
$c_{n,s}{\mathcal{L}}_{1}u(x)$ $=c_{n,s}{\rm p.v.}\int_{B_{1}(x)}\frac{u(x)-u(y)}{|x-y|^{n+2s}}\big(\!-2\ln|x-y|\big)\,dy+c_{n,s}\int_{\mathbb{R}^{n}\setminus B_{1}(x)}\frac{-2\ln|x-y|}{|x-y|^{n+2s}}\,dy\,u(x)$
$\qquad-c_{n,s}\int_{\mathbb{R}^{n}\setminus B_{1}(x)}\frac{u(y)}{|x-y|^{n+2s}}\big(\!-2\ln|x-y|\big)\,dy,$
where
$\Big|c_{n,s}{\rm p.v.}\int_{B_{1}(x)}\frac{u(x)-u(y)}{|x-y|^{n+2s}}\big(\!-2\ln|x-y|\big)\,dy\Big|$
$\leq$ $C_{0}c_{n,s}\int_{B_{1}(x)}\frac{1}{|x-y|^{n+2s-1}}\big(\!-2\ln|x-y|\big)\,dy$
$=$ $2C_{0}c_{n,s}\left|\mathbb{S}^{n-1}\right|(1-2s)^{-2}\to 0\qquad{\rm as}\ \ s\to 0^{+},$
and
$c_{n,s}\int_{\mathbb{R}^{n}\setminus B_{1}(x)}\frac{-2\ln|x-y|}{|x-y|^{n+2s}}\,dy\,u(x)=-2c_{n,s}\left|\mathbb{S}^{n-1}\right|(2s)^{-2}\,u(x).$
Moreover,
$\Big|c_{n,s}\int_{\mathbb{R}^{n}\setminus B_{1}(x)}\frac{u(y)}{|x-y|^{n+2s}}\big(\!-2\ln|x-y|\big)\,dy\Big|$
$\leq$ $c_{n,s}\|u\|_{L^{\infty}}\int_{B_{R}\setminus B_{1}(x)}\frac{2\ln|x-y|}{|x-y|^{n+2s}}\,dy$
$\leq$ $2c_{n,s}\left|\mathbb{S}^{n-1}\right|\|u\|_{L^{\infty}}\int_{1}^{R+1}\frac{\ln r}{r^{1+2s}}dr$
$=$ $2c_{n,s}\left|\mathbb{S}^{n-1}\right|\|u\|_{L^{\infty}}\left(-\frac{\ln(R+1)}{2s(R+1)^{2s}}+\frac{1-(R+1)^{-2s}}{4s^{2}}\right)\to 0\qquad{\rm as}\ \ s\to 0^{+}.$
Part 2: estimates for $b_{n,s}(-\Delta)^{s}u(x)$: Observe that
$b_{n,s}(-\Delta)^{s}u(x)$ $=b_{n,s}c_{n,s}\int_{B_{1}(x)}\frac{u(x)-u(y)}{|x-y|^{n+2s}}\,dy+b_{n,s}c_{n,s}\int_{\mathbb{R}^{n}\setminus B_{1}(x)}\frac{1}{|x-y|^{n+2s}}\,dy\,u(x)$
$\qquad-b_{n,s}c_{n,s}\int_{\mathbb{R}^{n}\setminus B_{1}(x)}\frac{u(y)}{|x-y|^{n+2s}}\,dy.$
Since for $\delta\in(0,1)$
$\Big|\int_{B_{\delta}(x)}\frac{u(x)-u(y)}{|x-y|^{n+2s}}\,dy\Big|$ $\leq C_{0}\int_{B_{\delta}(x)}\frac{1}{|x-y|^{n+2s-1}}\,dy$
$=C_{0}\left|\mathbb{S}^{n-1}\right|\int_{0}^{\delta}\frac{1}{r^{2s}}dr$
$=\frac{C_{0}}{1-2s}\left|\mathbb{S}^{n-1}\right|\delta^{1-2s}$
and
$\Big|\int_{B_{\delta}(x)}\frac{u(x)-u(y)}{|x-y|^{n}}\,dy\Big|$ $\leq C_{0}\int_{B_{\delta}(x)}\frac{1}{|x-y|^{n-1}}\,dy=C_{0}\left|\mathbb{S}^{n-1}\right|\delta$
then
$\Big|\int_{B_{1}(x)}\frac{u(x)-u(y)}{|x-y|^{n+2s}}\,dy-\int_{B_{1}(x)}\frac{u(x)-u(y)}{|x-y|^{n}}\,dy\Big|$
$\leq$ $\int_{B_{1}(x)\setminus B_{\delta}(x)}|u(x)-u(y)|\Big|\frac{1}{|x-y|^{n+2s}}-\frac{1}{|x-y|^{n}}\Big|dy+\frac{C_{0}}{1-2s}\left|\mathbb{S}^{n-1}\right|\delta^{1-2s}+C_{0}\left|\mathbb{S}^{n-1}\right|\delta$
$\leq$ $2\|u\|_{L^{\infty}}\int_{B_{1}(x)\setminus B_{\delta}(x)}\frac{1}{|x-y|^{n}}\Big(e^{2s\ln\frac{1}{|x-y|}}-1\Big)dy+\frac{C_{0}}{1-2s}\left|\mathbb{S}^{n-1}\right|\delta^{1-2s}+C_{0}\left|\mathbb{S}^{n-1}\right|\delta$
$\leq$ $\frac{4s}{\delta^{2s}}\ln\delta\|u\|_{L^{\infty}}\int_{\delta<|z|<1}\frac{1}{|z|^{n+1}}dz+\frac{C_{0}}{1-2s}\left|\mathbb{S}^{n-1}\right|\delta^{1-2s}+C_{0}\left|\mathbb{S}^{n-1}\right|\delta.$
Thus,
$\int_{B_{1}(x)}\frac{u(x)-u(y)}{|x-y|^{n+2s}}\,dy\to\int_{B_{1}(x)}\frac{u(x)-u(y)}{|x-y|^{n}}\,dy\quad\text{in}\quad L^{\infty}\quad s\rightarrow 0$
and
$b_{n,s}c_{n,s}\int_{B_{1}(x)}\frac{u(x)-u(y)}{|x-y|^{n+2s}}\,dy$ $\to\pi^{-\frac{n}{2}}\Gamma(\frac{n}{2})\int_{B_{1}(x)}\frac{u(x)-u(y)}{|x-y|^{n}}\,dy\quad\text{in}\quad L^{\infty}\quad s\rightarrow 0.$
Direct computations show that
$b_{n,s}c_{n,s}\int_{\mathbb{R}^{n}\setminus B_{1}(x)}\frac{1}{|x-y|^{n+2s}}\,dy\,u(x)$ $=b_{n,s}c_{n,s}\left|\mathbb{S}^{n-1}\right|\int_{1}^{\infty}\frac{1}{r^{2s+1}}dr\,u(x)$
$=b_{n,s}c_{n,s}\left|\mathbb{S}^{n-1}\right|(2s)^{-1}\,u(x).$
Part 3: estimates for $(-\Delta)^{s+\operatorname*{Log}}u(x)$:
$\quad c_{n,s}\int_{\mathbb{R}^{n}\setminus B_{1}(x)}\frac{-2\ln|x-y|}{|x-y|^{n+2s}}\,dy\,+b_{n,s}c_{n,s}\int_{\mathbb{R}^{n}\setminus B_{1}(x)}\frac{1}{|x-y|^{n+2s}}\,dy\,$
$=c_{n,s}\left|\mathbb{S}^{n-1}\right|\frac{1}{2s}\Big[b_{n,s}-\frac{1}{s}\Big]$
$=\left|\mathbb{S}^{n-1}\right|2^{2s-1}\pi^{-\frac{n}{2}}\frac{\Gamma(\frac{n+2s}{2})}{\Gamma(1-s)}\Big[\ln 4+\psi(1-s)+\psi(\frac{n+2s}{2})\Big]$
$\to\ln 4+\psi(1)+\psi(\frac{n}{2})=\rho_{n}\quad{\rm as}\ \ s\to 0^{+}.$
Thus, for $s\in(0,\frac{1}{4})$
$(-\Delta)^{s+\operatorname*{Log}}u(x)$ $=c_{n,s}{\mathcal{L}}_{1}u(x)+b_{n,s}(-\Delta)^{s}u(x)$
$=c_{n,s}\int_{B_{1}(x)}\frac{u(x)-u(y)}{|x-y|^{n+2s}}\big(\!-2\ln|x-y|\big)\,dy+b_{n,s}c_{n,s}\int_{B_{1}(x)}\frac{u(x)-u(y)}{|x-y|^{n+2s}}\,dy$
$+c_{n,s}\int_{\mathbb{R}^{n}\setminus B_{1}(x)}\frac{-2\ln|x-y|}{|x-y|^{n+2s}}\,dy\,u(x)+b_{n,s}c_{n,s}\int_{\mathbb{R}^{n}\setminus B_{1}(x)}\frac{1}{|x-y|^{n+2s}}\,dy\,u(x)$
$-c_{n,s}\int_{\mathbb{R}^{n}\setminus B_{1}(x)}\frac{u(y)}{|x-y|^{n+2s}}\big(\!-2\ln|x-y|\big)\,dy-b_{n,s}c_{n,s}\int_{\mathbb{R}^{n}\setminus B_{1}(x)}\frac{u(y)}{|x-y|^{n+2s}}\,dy$
$\to(-\Delta)^{\ln}u(x)\qquad{\rm as}\qquad\text{in}\qquad L^{\infty}\qquad s\to 0^{+}.$
The proof ends. ∎
2.2. Spectral Representation
In fact, the functional-calculus framework provides a clean and equivalent definition of the fractional–logarithmic Laplacian. We first recall the necessary notions.
Let $\Delta$ be the nonnegative self-adjoint Laplacian, which is an unbounded operator on $L^{2}(\mathbb{R}^{n})$. By the spectral theorem, there exists a spectral resolution $E(\cdot)$ on $0,\infty)$ such that
$\Delta=\int_{[0,\infty)}\lambda\,dE(\lambda).$
For each pair $u,v\in L^{2}(\mathbb{R}^{n})$, $E_{u,v}$ is a regular complex Borel measure of bounded variation on $0,\infty)$, supported on $\sigma(-\Delta)$, where
$E_{u,v}(B)\;=\;\bigl\langle E(B)\,u,\;v\bigr\rangle_{L^{2}(M,\mu)},\>\>E\left(\sigma\left(-\Delta\right)\right)=I\quad\text{where}\>\>B\subset[0,\infty)\text{ is Borel set}$
and satisfying
$\bigl|E_{u,v}\bigr|\bigl([0,\infty)\bigr)\;\leq\;\|u\|_{L^{2}(\mathbb{R}^{n})}\,\|v\|_{L^{2}(\mathbb{R}^{n})}.$
Moreover, $E_{u,u}$ is a positive measure with $E_{u,u}\bigl([0,\infty)\bigr)=\|u\|_{L^{2}(\mathbb{R}^{n})}.$ In fact, we have
$\operatorname{Dom}(-\Delta)=\Bigl\{u\in L^{2}(\mathbb{R}^{n}):\int_{[0,\infty)}\lambda^{2}dE_{u,u}(\lambda)<\infty\Bigr\},$
and
$\langle-\Delta u,v\rangle_{L^{2}}=\int_{[0,\infty)}\lambda\,dE_{u,v}(\lambda),\quad u\in\operatorname{Dom}(-\Delta),\ v\in L^{2}(\mathbb{R}^{n}).$
For every Borel function $\varphi:[0,\infty)\to\mathbb{R}$, one can define
$\varphi(-\Delta):=\int_{[0,\infty)}\varphi(\lambda)\,dE(\lambda),$
with
$\operatorname{Dom}(\varphi(-\Delta))=\Bigl\{u\in L^{2}(\mathbb{R}^{n}):\int_{[0,\infty)}|\varphi(\lambda)|^{2}dE_{u,u}(\lambda)<\infty\Bigr\}.$
The action rule
$\langle\varphi(-\Delta)u,v\rangle_{L^{2}}=\int_{[0,\infty)}\varphi(\lambda)\,dE_{u,v}(\lambda),\quad u\in\operatorname{Dom}(\varphi(-\Delta)),\ v\in L^{2}(\mathbb{R}^{n}).$
We next present a proposition which rigorously establishes that $(-\Delta)^{s+\operatorname*{Log}}$ coincides with the first-order derivative of the fractional Laplace $(-\Delta)^{t}$ with respect to $t$ at $t=s.$
Proof of Proposition 1.2:.
By spectral calculus, for $t$ close to $s$ we can write
$\frac{(-\Delta)^{t}-(-\Delta)^{s}}{t-s}\,u-(-\Delta)^{s+\operatorname*{Log}}u=\int_{[0,\infty)}h_{t,s}(\lambda)\,dE(\lambda)\,u,$
where
$h_{t,s}(\lambda):=\frac{\lambda^{t}-\lambda^{s}}{t-s}-\lambda^{s}\ln\lambda,\qquad\lambda>0,$
and we set $h_{t,s}(0):=0$.
Hence, using the spectral measure $E_{u,u}$,
(2.2) $\Bigl\|\frac{(-\Delta)^{t}-(-\Delta)^{s}}{t-s}\,u-(-\Delta)^{s+\operatorname*{Log}}u\Bigr\|_{L^{2}}^{2}=\int_{[0,\infty)}|h_{t,s}(\lambda)|^{2}\,dE_{u,u}(\lambda).$
For each fixed $\lambda>0$, we have $h_{t,s}(\lambda)\to 0$ as $t\to s$, since $\partial_{t}(\lambda^{t})=\lambda^{t}\ln\lambda$.
It remains to control $h_{t,s}$ uniformly in $\lambda$ for $t$ near $s$. Fix $\delta\in(0,\min\left\{s,s_{0}-s\right\})$ and restrict to $|t-s|<\delta$. Consider the smooth function $F(\tau):=\lambda^{\tau}$ on $\tau\in\mathbb{R}$. By Taylor’s theorem with remainder, there exists $\theta=\theta(\lambda,t,s)$ between $s$ and $t$ such that
$\lambda^{t}=\lambda^{s}+(t-s)\lambda^{s}\ln\lambda+\frac{(t-s)^{2}}{2}\lambda^{\theta}(\ln\lambda)^{2}.$
Therefore,
$h_{t,s}(\lambda)=\frac{(t-s)}{2}\lambda^{\theta}(\ln\lambda)^{2},$
and hence, for $|t-s|<\delta$,
(2.3) $|h_{t,s}(\lambda)|^{2}\leq C\,|t-s|^{2}\,(\ln\lambda)^{4}\Bigl(\lambda^{2(s-\delta)}\mathbf{1}_{\{\lambda\leq 1\}}+\lambda^{2(s+\delta)}\mathbf{1}_{\{\lambda\geq 1\}}\Bigr),$
with a constant $C>0$ independent of $\lambda,t$.
Since $s-\delta>0$, the function $\lambda\mapsto(\ln\lambda)^{4}\lambda^{2(s-\delta)}$ is bounded on $(0,1]$. Hence
$\int_{(0,1]}(\ln\lambda)^{4}\lambda^{2(s-\delta)}\,dE_{u,u}(\lambda)\leq C\int_{(0,1]}dE_{u,u}(\lambda)\leq C\|u\|_{L^{2}}^{2}<\infty.$
Since $\delta<s_{0}-s$ and $\ln\lambda$ grows slower than any power of $\lambda$, there exists $C_{\delta}>0$ such that for all $\lambda\geq 1$,
$(\ln\lambda)^{4}\,\lambda^{2(s+\delta)}\leq C_{\delta}\,\lambda^{2s_{0}}.$
Therefore, using $u\in H^{2s_{0}}(\mathbb{R}^{n})$,
$\int_{[1,\infty)}(\ln\lambda)^{4}\lambda^{2(s+\delta)}\,dE_{u,u}(\lambda)\leq C_{\delta}\int_{[1,\infty)}\lambda^{2s_{0}}\,dE_{u,u}(\lambda)<\infty.$
Combining these two bounds with (2.3), we obtain from (2.2) that
$\int_{[0,\infty)}|h_{t,s}(\lambda)|^{2}\,dE_{u,u}(\lambda)\leq C\,|t-s|^{2},$
where $C$ is independent of $t$ near $s$. Hence the left-hand side tends to 0 as $t\to s$. ∎
In what follows, we relate the spectral measure of $\Delta$ on $\mathbb{R}^{n}$ to the Fourier transform, and thus provide fourier-multiplier characterizations of the fractional–logarithmic operator and the associated space.
In the Euclidean setting, $\Delta$ is unitarily diagonalized by the Fourier transform. More precisely, for $f\in L^{2}(\mathbb{R}^{n})$, the spectral measure of $\Delta$ satisfies
$E_{f,f}(B)=\int_{\{|\xi|^{2}\in B\}}|\widehat{f}(\xi)|^{2}\,d\xi,\qquad B\subset[0,\infty)\ \text{Borel}.$
Writing
$F(\lambda):=E_{f,f}\bigl((0,\lambda]\bigr)=\int_{|\xi|\leq\sqrt{\lambda}}|\widehat{f}(\xi)|^{2}\,d\xi$
and differentiating in $\lambda$ in polar coordinates $\xi=r\omega$ yields
$dE_{f,f}(\lambda)=F^{\prime}(\lambda)\,d\lambda=\frac{1}{2}\,\lambda^{\frac{n}{2}-1}\int_{\mathbb{S}^{n-1}}\bigl|\widehat{f}(\sqrt{\lambda}\,\omega)\bigr|^{2}\,d\omega\,d\lambda.$
Lemma 2.1.
Let $u,v\in L^{2}(\mathbb{R}^{n})$. Then for every Borel set $B\subset[0,\infty)$,
$E_{u,v}(B):=\langle E(B)u,v\rangle_{L^{2}}=\int_{\{|\xi|^{2}\in B\}}\widehat{u}(\xi)\,\overline{\widehat{v}(\xi)}\,d\xi.$
In particular, writing
$F_{u,v}(\lambda):=E_{u,v}((0,\lambda])=\int_{|\xi|\leq\sqrt{\lambda}}\widehat{u}(\xi)\,\overline{\widehat{v}(\xi)}\,d\xi,$
one has that $E_{u,v}$ is absolutely continuous on $(0,\infty)$ and
(2.4) $dE_{u,v}(\lambda)=F^{\prime}_{u,v}(\lambda)\,d\lambda=\frac{1}{2}\,\lambda^{\frac{n}{2}-1}\int_{\mathbb{S}^{n-1}}\widehat{u}(\sqrt{\lambda}\,\omega)\,\overline{\widehat{v}(\sqrt{\lambda}\,\omega)}\,d\omega\,d\lambda.$
Proof.
Since $\mathcal{F}$ is unitary and diagonalizes $\Delta$ on $\mathbb{R}^{n}$, the spectral projection $E(B)$ acts as a Fourier multiplier:
$\widehat{E(B)f}(\xi)=\mathbf{1}_{B}(|\xi|^{2})\,\widehat{f}(\xi),\qquad f\in\mathscr{S}(\mathbb{R}^{n}).$
Therefore, by Plancherel,
$E_{u,v}(B)=\langle E(B)u,v\rangle_{L^{2}}=\int_{\mathbb{R}^{n}}\mathbf{1}_{B}(|\xi|^{2})\,\widehat{u}(\xi)\,\overline{\widehat{v}(\xi)}\,d\xi=\int_{\{|\xi|^{2}\in B\}}\widehat{u}(\xi)\,\overline{\widehat{v}(\xi)}\,d\xi.$
For $\lambda>0$,
$F_{u,v}(\lambda)=E_{u,v}((0,\lambda])=\int_{|\xi|\leq\sqrt{\lambda}}\widehat{u}(\xi)\,\overline{\widehat{v}(\xi)}\,d\xi.$
Passing to polar coordinates $\xi=r\omega$,
$F_{u,v}(\lambda)=\int_{0}^{\sqrt{\lambda}}\int_{\mathbb{S}^{n-1}}\widehat{u}(r\omega)\,\overline{\widehat{v}(r\omega)}\,r^{n-1}\,d\omega\,dr.$
Since $u,v\in\mathscr{S}(\mathbb{R}^{n})$, the integrand is smooth and integrable; thus we may differentiate with respect to $\lambda$:
$F^{\prime}_{u,v}(\lambda)=\frac{1}{2\sqrt{\lambda}}\int_{\mathbb{S}^{n-1}}\widehat{u}(\sqrt{\lambda}\,\omega)\,\overline{\widehat{v}(\sqrt{\lambda}\,\omega)}\,(\sqrt{\lambda})^{\,n-1}\,d\omega=\frac{1}{2}\,\lambda^{\frac{n}{2}-1}\int_{\mathbb{S}^{n-1}}\widehat{u}(\sqrt{\lambda}\,\omega)\,\overline{\widehat{v}(\sqrt{\lambda}\,\omega)}\,d\omega,$
which gives (2.4). ∎
Lemma 2.2.
For every $s\in(0,1)$ and $u\in H^{s+\operatorname*{Log}}(\mathbb{R}^{n}),\ v\in L^{2}(\mathbb{R}^{n})$
(2.5) $\int_{[0,\infty)}\lambda^{s}\ln\lambda\,dE_{u,v}(\lambda)=\int_{\mathbb{R}^{n}}|\xi|^{2s}\ln\bigl(|\xi|^{2}\bigr)\,\widehat{u}(\xi)\,\overline{\widehat{v}(\xi)}\,d\xi.$
Proof.
Using (2.4),
$\int_{[0,\infty)}\lambda^{s}\ln\lambda\,dE_{u,v}(\lambda)$ $=\int_{0}^{\infty}\lambda^{s}\ln\lambda\,F^{\prime}_{u,v}(\lambda)\,d\lambda$
$=\frac{1}{2}\int_{0}^{\infty}\lambda^{\frac{n}{2}-1+s}\ln\lambda\int_{\mathbb{S}^{n-1}}\widehat{u}(\sqrt{\lambda}\,\omega)\,\overline{\widehat{v}(\sqrt{\lambda}\,\omega)}\,d\omega\,d\lambda.$
Now set $\lambda=r^{2}$, then
$\int_{[0,\infty)}\lambda^{s}\ln\lambda\,dE_{u,v}(\lambda)$ $=\frac{1}{2}\int_{0}^{\infty}\!\!\int_{\mathbb{S}^{n-1}}r^{n-2+2s}\ln(r^{2})\,\widehat{u}(r\omega)\,\overline{\widehat{v}(r\omega)}\,d\omega\,(2r\,dr)$
$=\int_{0}^{\infty}\!\!\int_{\mathbb{S}^{n-1}}r^{n-1+2s}\ln(r^{2})\,\widehat{u}(r\omega)\,\overline{\widehat{v}(r\omega)}\,d\omega\,dr$
$=\int_{\mathbb{R}^{n}}|\xi|^{2s}\ln\bigl(|\xi|^{2}\bigr)\,\widehat{u}(\xi)\,\overline{\widehat{v}(\xi)}\,d\xi,$
which is exactly (2.5). ∎
Proof of Theorem 1.2:.
(i) We now change variables $\lambda=r^{2}$, then
$\int_{0}^{\infty}\lambda^{s}\ln^{2}\lambda\,dE_{f,f}(\lambda)$ $=\frac{1}{2}\int_{0}^{\infty}\lambda^{\frac{n}{2}-1+s}\ln^{2}\lambda\int_{\mathbb{S}^{n-1}}\bigl|\widehat{f}(\sqrt{\lambda}\,\omega)\bigr|^{2}\,d\omega\,d\lambda$
$=\frac{1}{2}\int_{0}^{\infty}\!\!\int_{\mathbb{S}^{n-1}}r^{n-2+2s}\ln^{2}(r^{2})\bigl|\widehat{f}(r\omega)\bigr|^{2}\,d\omega\,2r\,dr$
$=\int_{0}^{\infty}\!\!\int_{\mathbb{S}^{n-1}}r^{n-1+2s}\ln^{2}(r^{2})\bigl|\widehat{f}(r\omega)\bigr|^{2}\,d\omega\,dr$
$=\int_{\mathbb{R}^{n}}|\xi|^{2s}\ln^{2}\bigl(|\xi|^{2}\bigr)\,\bigl|\widehat{f}(\xi)\bigr|^{2}\,d\xi.$
(ii) Let $\varphi\in L^{2}(\mathbb{R}^{n})$. By the definition of the functional calculus,
$\langle(-\Delta)^{s+\operatorname*{Log}}u,\varphi\rangle_{L^{2}}=\int_{[0,\infty)}\lambda^{s}\ln\lambda\,dE_{u,\varphi}(\lambda).$
Applying Lemma 2.2 with $v=\varphi$ gives
$\langle(-\Delta)^{s+\operatorname*{Log}}u,\varphi\rangle_{L^{2}}=\int_{\mathbb{R}^{n}}|\xi|^{2s}\ln(|\xi|^{2})\,\widehat{u}(\xi)\,\overline{\widehat{\varphi}(\xi)}\,d\xi.$
By Plancherel, the right-hand side equals
$\Big\langle{\mathcal{F}}^{-1}\!\big(|\xi|^{2s}\ln(|\xi|^{2})\,\widehat{u}\big),\,\varphi\Big\rangle_{L^{2}}.$
Since $\mathscr{S}(\mathbb{R}^{n})$ is dense in $L^{2}(\mathbb{R}^{n})$, we conclude that
$(-\Delta)^{s+\operatorname*{Log}}u={\mathcal{F}}^{-1}\!\big(|\xi|^{2s}\ln(|\xi|^{2})\,\widehat{u}\big)\quad\text{in }L^{2}(\mathbb{R}^{n}),$
which is equivalent to (1.9). The characterization of the domain follows immediately from Plancherel:
$(-\Delta)^{s+\operatorname*{Log}}u\in L^{2}\iff|\xi|^{2s}\ln(|\xi|^{2})\,\widehat{u}(\xi)\in L^{2}.$
2.3. Extension Problem
In this subsection, we derive the extension characterization of the fractional-logarithmic operator and identify the associated Dirichlet-to-Neumann-type boundary formula. We first establish the boundary trace of $v_{s}$.
Lemma 2.3.
Let $v_{s}$ be defined in (1.11), then
$\lim_{t\to 0^{+}}v_{s}(x,t)=b_{1}u(x),$
where
$b_{1}:=p_{n,s}\int_{\mathbb{R}^{n}}\frac{-\ln(|z|^{2}+1)}{(|z|^{2}+1)^{\frac{n+2s}{2}}}dz=-\frac{\partial_{s}p_{n,s}}{p_{n,s}}.$
Proof.
Direct computation shows that
$v_{s}(x,t)$ $=p_{n,s}\ t^{2s}\int_{\mathbb{R}^{n}}\frac{2\ln t-\ln(|x-y|^{2}+t^{2})}{(|x-y|^{2}+t^{2})^{\frac{n+2s}{2}}}u(y)dy$
$=p_{n,s}\int_{\mathbb{R}^{n}}\frac{-\ln(z^{2}+1)}{(z^{2}+1)^{\frac{n+2s}{2}}}u(x+tz)dz\to b_{1}u(x)\quad{\rm as}\ \,t\to 0^{+},$
where
$b_{1}=p_{n,s}\int_{\mathbb{R}^{n}}\frac{-\ln(|z|^{2}+1)}{(|z|^{2}+1)^{\frac{n+2s}{2}}}dz.$
Set
$I(s):=\int_{\mathbb{R}^{n}}\frac{1}{(1+|z|^{2})^{\frac{n+2s}{2}}}\,dz,\qquad\text{so that}\qquad p_{n,s}=\frac{1}{I(s)}.$
For $s\in(0,1)$ the integrand is smooth in $s$ and satisfies
$\partial_{s}\Big((1+|z|^{2})^{-\frac{n+2s}{2}}\Big)=-(\ln(1+|z|^{2}))\,(1+|z|^{2})^{-\frac{n+2s}{2}}.$
Moreover, since for large $|z|$ we have
$|\ln(1+|z|^{2})|\,(1+|z|^{2})^{-\frac{n+2s}{2}}\lesssim(1+|z|^{2})^{-\frac{n+s}{2}},$
the derivative is integrable, hence we may differentiate under the integral sign to obtain
$I^{\prime}(s)=\int_{\mathbb{R}^{n}}\frac{-\ln(1+|z|^{2})}{(1+|z|^{2})^{\frac{n+2s}{2}}}\,dz.$
Therefore, by the definition of $b_{1}$,
$b_{1}=p_{n,s}\,I^{\prime}(s)=\frac{I^{\prime}(s)}{I(s)}.$
On the other hand, since $p_{n,s}=I(s)^{-1}$,
$\frac{\partial_{s}p_{n,s}}{p_{n,s}}=\partial_{s}\big(\ln p_{n,s}\big)=\partial_{s}\big(-\ln I(s)\big)=-\frac{I^{\prime}(s)}{I(s)}=-b_{1}.$
This proves the claim, we complete the proof. ∎
We now turn to the corresponding extension problem and complete the proof.
Proof of Theorem 1.3.
Note that
(2.6) $\partial_{s}w_{s}=\frac{\partial_{s}p_{n,s}}{p_{n,s}}w_{s}+v_{s}$
and
$\lim_{t\to 0^{+}}\partial_{s}w_{s}(x,t)$ $=\lim_{t\to 0^{+}}\frac{\partial_{s}p_{n,s}}{p_{n,s}}w_{s}(x,t)+\lim_{t\to 0^{+}}v_{s}(x,t)=\frac{\partial_{s}p_{n,s}}{p_{n,s}}u(x)+b_{1}u(x)=0.$
Observe that
(2.7) $v_{s}(x,t)=t^{2s}(2\ln t)w_{s}(x,t)+t^{2s}\bar{v}_{s}(x,t),$
where
$\bar{v}_{s}=p_{n,s}\int_{\mathbb{R}^{n}}\frac{-\ln(|x-y|^{2}+t^{2})}{(|x-y|^{2}+t^{2})^{\frac{n+2s}{2}}}u(y)dy.$
Moreover, we see that
$t^{-2s}\big(v_{s}(x,t)-v_{s}(x,0)\big)$ $=p_{n,s}\int_{\mathbb{R}^{n}}\frac{2\ln t-\ln(|x-y|^{2}+t^{2})}{(|x-y|^{2}+t^{2})^{\frac{n+2s}{2}}}\big(u(y)-u(x)\big)dy$
$=p_{n,s}(2\ln t)\int_{\mathbb{R}^{n}}\frac{1}{(|x-y|^{2}+t^{2})^{\frac{n+2s}{2}}}\big(u(y)-u(x)\big)dy+$
$\quad\ p_{n,s}\int_{\mathbb{R}^{n}}\frac{-\ln(|x-y|^{2}+t^{2})}{(|x-y|^{2}+t^{2})^{\frac{n+2s}{2}}}\big(u(y)-u(x)\big)dy$
$=(2\ln t)\,t^{-2s}\big(w_{s}(x,t)-w_{s}(x,0)\big)+$
$\quad\ p_{n,s}\int_{\mathbb{R}^{n}}\frac{-\ln(|x-y|^{2}+t^{2})}{(|x-y|^{2}+t^{2})^{\frac{n+2s}{2}}}\big(u(y)-u(x)\big)\,dy,$
and
$\lim_{t\to 0^{+}}{\bf d}_{s}t^{-2s}\Big(w_{s}(x,t)-w_{s}(x,0)\Big)$
$=$ ${\bf d}_{s}p_{n,s}\lim_{t\to 0^{+}}t^{-2s}\Big(\int_{\mathbb{R}^{n}}\frac{t^{2s}u(\tilde{x})}{(|x-\tilde{x}|^{2}+t^{2})^{\frac{n+2s}{2}}}d\tilde{x}-\int_{\mathbb{R}^{n}}\frac{t^{2s}}{(|z|^{2}+t^{2})^{\frac{n+2s}{2}}}dzu(x)\Big)$
$=$ ${\bf d}_{s}p_{n,s}\lim_{t\to 0^{+}}\int_{\mathbb{R}^{n}}\frac{u(\tilde{x})-u(x)}{(|x-\tilde{x}|^{2}+t^{2})^{\frac{n+2s}{2}}}d\tilde{x}$
$=$ $-(-\Delta)^{s}u(x).$
It is known that
$\lim_{t\to 0^{+}}w_{s}(x,t)=u(x)$
and
$\lim_{t\to 0^{+}}{\bf d}_{s}t^{1-2s}\partial_{t}w_{s}(x,t)=-(-\Delta)^{s}u(x).$
Then we obtain that
$\partial_{s}(-\Delta)^{s}u(x)$ $=-\lim_{t\to 0^{+}}\big[\partial_{s}\big({\bf d}_{s}t^{-2s}\big(w_{s}(x,t)-w_{s}(x,0)\big)\big)\big]$
$=-\lim_{t\to 0^{+}}\Big[\left\{{\bf d}_{s}^{\prime}+{\bf d}_{s}(-2\ln t)+{\bf d}_{s}\frac{p_{n,s}^{\prime}}{p_{n,s}}\right\}t^{-2s}\big(w_{s}(x,t)-w_{s}(x,0)\big)+$
$\qquad{\bf d}_{s}t^{-2s}\big(v_{s}(x,t)-v_{s}(x,0)\big)\Big]$
$=-\lim_{t\to 0^{+}}\Big[\big({\bf d}_{s}^{\prime}+{\bf d}_{s}\frac{p_{n,s}^{\prime}}{p_{n,s}}\big)t^{-2s}\big(w_{s}(x,t)-w_{s}(x,0)\big)+$
$\qquad{\bf d}_{s}p_{n,s}\int_{\mathbb{R}^{n}}\frac{-\ln(|x-y|^{2}+t^{2})}{(|x-y|^{2}+t^{2})^{\frac{n+2s}{2}}}\big(u(y)-u(x)\big)dy\Big]$
$=k_{s}\ (-\Delta)^{s}u+c_{n,s}{\mathcal{L}}_{1}u,$
where
$k_{s}=\frac{{\bf d}_{s}^{\prime}}{{\bf d}_{s}}+\frac{\partial_{s}p_{n,s}}{p_{n,s}}.$
Direct computation shows that
$\frac{{\bf d}_{s}^{\prime}}{{\bf d}_{s}}=\ln 4+\psi(s)+\psi(1-s)+\frac{1}{s}$
and
$\frac{\partial_{s}p_{n,s}}{p_{n,s}}=-\psi(s)+\psi(\frac{n+2s}{2}).$
Thus, we obtain that
$k_{s}=\ln 4+\frac{1}{s}+\psi(1-s)+\psi(\frac{n+2s}{2})=b_{n,s},$
then we have
$\partial_{s}(-\Delta)^{s}u(x)=(-\Delta)^{s+\operatorname*{Log}}u(x).$
Since $w_{s}$ solves
${\rm div}\,_{(x,t)}\!\big(t^{1-2s}\nabla_{(x,t)}w_{s}\big)=0\qquad\text{in }\mathbb{R}^{n+1}_{+},$
differentiating with respect to $s$ gives
$0$ $=\partial_{s}\,{\rm div}\,_{(x,t)}\!\big(t^{1-2s}\nabla_{(x,t)}w_{s}\big)$
$=\,{\rm div}\,_{(x,t)}\!\big(\partial_{s}(t^{1-2s})\,\nabla_{(x,t)}w_{s}\big)+\,{\rm div}\,_{(x,t)}\!\big(t^{1-2s}\nabla_{(x,t)}(\partial_{s}w_{s})\big)$
$=\,{\rm div}\,_{(x,t)}\!\big(t^{1-2s}(-2\ln t)\nabla_{(x,t)}w_{s}\big)+\,{\rm div}\,_{(x,t)}\!\big(t^{1-2s}\nabla_{(x,t)}(\partial_{s}w_{s})\big).$
Using the product rule and ${\rm div}\,_{(x,t)}(t^{1-2s}\nabla_{(x,t)}w_{s})=0$, we further obtain
${\rm div}\,_{(x,t)}\!\big(t^{1-2s}(-2\ln t)\nabla_{(x,t)}w_{s}\big)=\nabla_{(x,t)}(-2\ln t)\cdot t^{1-2s}\nabla_{(x,t)}w_{s}=-2t^{-2s}\partial_{t}w_{s}.$
Finally, by (2.6) we know $v_{s}=\partial_{s}w_{s}-\frac{p^{\prime}_{n,s}}{p_{n,s}}\,w_{s}$ and ${\rm div}\,_{(x,t)}(t^{1-2s}\nabla_{(x,t)}w_{s})=0$, it follows that
${\rm div}\,_{(x,t)}\!\big(t^{1-2s}\nabla_{(x,t)}(\partial_{s}w_{s})\big)=\,{\rm div}\,_{(x,t)}\!\big(t^{1-2s}\nabla_{(x,t)}v_{s}\big).$
Consequently,
${\rm div}\,_{(x,t)}\!\big(t^{1-2s}\nabla_{(x,t)}v_{s}\big)=2t^{-2s}\partial_{t}w_{s}\qquad\text{in }\mathbb{R}^{n+1}_{+}.$
As a consequence, $v_{s}$ verifies the equation (1.13) and
$2t^{-2s}\partial_{t}w_{s}(x,t)=4st^{-1-2s}w_{s}(x,t)+2(n+2s)p_{n,s}t\int_{\mathbb{R}^{n}}\frac{u(y)}{(|x-y|^{2}+t^{2})^{\frac{n+2s}{2}+1}}dy,$
which ends the proof. ∎
3. Poisson Problem
In this section, we study the Poisson problem associated with the fractional-logarithmic operator, focusing on existence and regularity of weak solutions. The existence theory is based mainly on the Lax–Milgram framework, while the regularity part relies on a Moser iteration argument. To carry this out, we first establish the required structural properties of the corresponding function spaces and energy forms.
3.1. Properties of Functional Spaces
We first clarify the relationship between the spectral fractional-logarithmic space and the energy fractional-logarithmic space, which helps establish a transparent link between functional calculus and the Fourier-transform approach.
Proof of Proposition 1.3:.
Set z=xyz=x-y and arguing as in the standard fractional Sobolev case, see [11, Proposition 3.4] one gets
(3.1) $\iint_{\mathbb{R}^{n}\times\mathbb{R}^{n}}\bigl(u(x)-u(y)\bigr)^{2}\,{\bf k}_{s+\operatorname*{Log},+}(x-y)\,dx\,dy=\int_{\mathbb{R}^{n}}m(\xi)\,|\widehat{u}(\xi)|^{2}\,d\xi,$
where
$m(\xi):=2\int_{\mathbb{R}^{n}}\bigl(1-\cos(\xi\cdot z)\bigr)\,{\bf k}_{s+\operatorname*{Log},+}(z)\,dz.$
Since ${\bf k}_{s+\operatorname*{Log},+}$ is radial and nonnegative, $m(\xi)=m(|\xi|)\geq 0$.
Write $r:=|\xi|$. Using $1-\cos(\xi\cdot z)\leq\min\{2,\,|\xi\cdot z|^{2}\}\leq\min\{2,\,r^{2}|z|^{2}\}$ and that ${\bf k}_{s+\operatorname*{Log},+}$ is supported in $\{|z|<1\}$, we split
$m(\xi)\leq 2\!\!\int_{|z|<1/r}\!\!r^{2}|z|^{2}\,{\bf k}_{s+\operatorname*{Log},+}(z)\,dz+4\!\!\int_{1/r\leq|z|<1}\!\!{\bf k}_{s+\operatorname*{Log},+}(z)\,dz.$
A direct radial computation gives, for $r\geq 2$,
$\int_{|z|<1/r}r^{2}|z|^{2}\,|z|^{-n-2s}(-\ln|z|)\,dz\ \lesssim\ r^{2s}\ln r,\ \int_{1/r\leq|z|<1}|z|^{-n-2s}(-\ln|z|)\,dz\ \lesssim\ r^{2s}\ln r,$
hence $m(\xi)\lesssim r^{2s}\ln r$ for $r\geq 2$. For $0<r\leq 2$ we have $m(\xi)\lesssim 1+r^{2}$, since $1-\cos(\xi\cdot z)\leq Cr^{2}|z|^{2}$ for $|z|<1$ and
$\int_{|z|<1}|z|^{2}\,|z|^{-n-2s}(-\ln|z|)\,dz<\infty\qquad s\in(0,1).$
Combining both regimes yields the uniform bound
(3.2) $m(\xi)\ \leq\ C\bigl(1+|\xi|^{2s}(1+|\ln|\xi||)\bigr)\qquad\text{for all }\xi\in\mathbb{R}^{n}.$
Therefore,
$\mathcal{H}^{s+\operatorname*{Log}}(\mathbb{R}^{n}):=\Bigl\{u\in L^{2}(\mathbb{R}^{n}):\ \int_{\mathbb{R}^{n}}|\xi|^{2s}\ln(|\xi|^{2})|\widehat{u}(\xi)|^{2}\,d\xi<\infty\Bigr\},$
proving the continuous embedding $H^{s+\operatorname*{Log}}(\mathbb{R}^{n})\hookrightarrow\mathcal{H}^{s+\operatorname*{Log}}(\mathbb{R}^{n})$.
The form norm only controls one logarithmic power in the multiplier, whereas $H^{s+\operatorname*{Log}}$ requires square integrability with weight $\ln^{2}$. Thus one can construct $u$ with $\int|\xi|^{2s}|\ln|\xi||\,|\widehat{u}|^{2}<\infty$ but $\int|\xi|^{4s}\ln^{2}(|\xi|^{2})\,|\widehat{u}|^{2}=\infty$, showing that the reverse inclusion fails in general. ∎
Proposition 3.1.
The space $\mathcal{H}^{s+\operatorname*{Log}}(\mathbb{R}^{n})$, endowed with the inner product
$\langle u,v\rangle_{\mathcal{H}^{s+\operatorname*{Log}}}:=\int_{\mathbb{R}^{n}}u(x)v(x)\,dx+\iint_{\mathbb{R}^{n}\times\mathbb{R}^{n}}\bigl(u(x)-u(y)\bigr)\bigl(v(x)-v(y)\bigr)\,{\bf k}_{s+\operatorname*{Log},+}(x-y)\,dx\,dy,$
is a Hilbert space. In particular,
$\|u\|_{\mathcal{H}^{s+\operatorname*{Log}}}^{\,2}=\langle u,u\rangle_{\mathcal{H}^{s+\operatorname*{Log}}},\qquad u\in\mathcal{H}^{s+\operatorname*{Log}}(\mathbb{R}^{n}).$
Proof.
For $u,v\in\mathcal{H}^{s+\operatorname*{Log}}(\mathbb{R}^{n})$, the bilinear form is well-defined. Indeed, by Cauchy–Schwarz,
$\left|\iint(u(x)-u(y))(v(x)-v(y))\,{\bf k}_{s+\operatorname*{Log},+}(x-y)\,dx\,dy\right|\leq\,[u]_{s+\operatorname*{Log},+}\,[v]_{s+\operatorname*{Log},+}<\infty.$
If $\langle u,u\rangle_{\mathcal{H}^{s+\operatorname*{Log}}}=0$, then $\|u\|_{L^{2}(\mathbb{R}^{n})}^{\,2}=0$, hence $u=0$ a.e. in $\mathbb{R}^{n}$. Therefore, it is an inner product.
Let $(u_{k})_{k\in\mathbb{N}}$ be a Cauchy sequence in $\mathcal{H}^{s+\operatorname*{Log}}(\mathbb{R}^{n})$. Then it is Cauchy in $L^{2}(\mathbb{R}^{n})$, so there exists $u\in L^{2}(\mathbb{R}^{n})$ such that $u_{k}\to u$ in $L^{2}$. Moreover, since $(u_{k})$ is Cauchy, the sequence is bounded in the $\mathcal{H}^{s+\operatorname*{Log}}$-norm; thus,
$\sup_{k}[u_{k}]_{s+\operatorname*{Log},+}<\infty.$
Passing to a subsequence (not relabeled), we may assume $u_{k}(x)\to u(x)$ for a.e. $x\in\mathbb{R}^{n}$. By Fatou’s lemma and the nonnegativity of the kernel ${\bf k}_{s+\operatorname*{Log},+}$, we obtain
$\iint_{\mathbb{R}^{n}\times\mathbb{R}^{n}}\bigl(u(x)-u(y)\bigr)^{2}\,{\bf k}_{s+\operatorname*{Log},+}(x-y)\,dx\,dy$
$\leq$ $\liminf_{k\to\infty}\iint_{\mathbb{R}^{n}\times\mathbb{R}^{n}}\bigl(u_{k}(x)-u_{k}(y)\bigr)^{2}\,{\bf k}_{s+\operatorname*{Log},+}(x-y)\,dx\,dy,$
which shows that $u]_{s+\operatorname*{Log},+}<\infty$, i.e. $u\in\mathcal{H}^{s+\operatorname*{Log}}(\mathbb{R}^{n})$.
Finally, we show that $u_{k}\to u$ in $\mathcal{H}^{s+\operatorname*{Log}}$. Fix $\varepsilon>0$. Since $(u_{k})$ is Cauchy in $\mathcal{H}^{s+\operatorname*{Log}}$, there exists $N$ such that for all $k,m\geq N$,
$\|u_{k}-u_{m}\|_{L^{2}}^{2}+[u_{k}-u_{m}]_{s+\operatorname*{Log},+}^{2}<\varepsilon.$
Letting $m\to\infty$ and using $u_{m}\to u$ in $L^{2}$ we get $\|u_{k}-u\|_{L^{2}}^{2}\leq\varepsilon$ for all $k\geq N$. Moreover, by Fatou’s lemma, we obtain
$u_{k}-u]_{s+\operatorname*{Log},+}^{2}\leq\liminf_{m\to\infty}[u_{k}-u_{m}]_{s+\operatorname*{Log},+}^{2}\leq\varepsilon,\qquad k\geq N.$
Hence $\|u_{k}-u\|_{\mathcal{H}^{s+\operatorname*{Log}}}^{2}=\|u_{k}-u\|_{L^{2}}^{2}+[u_{k}-u]_{s+\operatorname*{Log},+}^{2}\leq 2\varepsilon$ for all $k\geq N$. This proves that $\mathcal{H}^{s+\operatorname*{Log}}(\mathbb{R}^{n})$ is complete, and therefore a Hilbert space. ∎
We now prove the classical Poincaré inequality associated with the fractional-logarithmic operator.
Proof of Proposition 1.4:.
Set
$K(z):={\bf k}_{s+\operatorname*{Log},+}(z)=c_{n,s}|z|^{-n-2s}(-\ln|z|)_{+},\quad z\in\mathbb{R}^{n}\setminus\{0\},$
then
$K(z)\ \geq\ c_{n,s}|z|^{-n-2s}\,\mathbf{1}_{\{|z|\leq e^{-1}\}}.$
Moreover, $K\geq 0$ and $K$ is integrable at infinity. Therefore, similar to the proof of [3, Lemma 1.28], we obtain
(3.3) $\|u\|_{L^{2_{s}^{*}}(\Omega)}^{\,2}=\|u\|_{L^{2_{s}^{*}}(\mathbb{R}^{n})}^{\,2}\ \leq\ C_{1}\iint_{\mathbb{R}^{n}\times\mathbb{R}^{n}}\bigl(u(x)-u(y)\bigr)^{2}K(x-y)\,dx\,dy=C_{1}[u]_{s+\operatorname*{Log},+}^{2},$
for some $C_{1}=C_{1}(n,s)>0$.
Finally, since $\Omega$ is bounded, Hölder’s inequality gives
$\|u\|_{L^{2}(\Omega)}^{2}\leq|\Omega|^{1-\frac{2}{2_{s}^{*}}}\,\|u\|_{L^{2_{s}^{*}}(\Omega)}^{2}=|\Omega|^{\frac{2s}{n}}\,\|u\|_{L^{2_{s}^{*}}(\Omega)}^{2}.$
Combining this with (3.3) we obtain (1.14) with $C:=|\Omega|^{\frac{2s}{n}}C_{1}$. ∎
We now prove the property of energy-forms .
Proof of Proposition 1.5:.
(i) Note that
$\mathcal{E}_{-}(u,u)$ $=c_{n,s}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\big(u(x)-u(y)\big)^{2}}{|x-y|^{n+2s}}(-\ln|x-y|\big)_{-}dxdy$
$\leq 4c_{n,s}\int_{\mathbb{R}^{n}}u(x)^{2}\int_{\mathbb{R}^{n}}\frac{(\ln|x-y|\big)_{+}}{|x-y|^{n+2s}}dydx$
$=4c_{n,s}\left|\mathbb{S}^{n-1}\right|\|u\|_{L^{2}(\Omega)}^{2}\int_{1}^{\infty}\frac{\ln r}{r^{2s+1}}dr$
$=\frac{1}{s^{2}}c_{n,s}\left|\mathbb{S}^{n-1}\right|\|u\|_{L^{2}(\Omega)}^{2},$
by using the fact
$\int_{1}^{\infty}\frac{\ln r}{r^{2s+1}}dr=\frac{1}{4s^{2}}.$
(ii) Recall that for $u\in\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)$ one has $u\equiv 0$ a.e. in $\Omega^{c}$ and
$\mathcal{E}_{+}(u,u)-\mathcal{E}_{-}(u,u)$ $=c_{n,s}\iint_{\mathbb{R}^{n}\times\mathbb{R}^{n}}\frac{(u(x)-u(y))^{2}}{|x-y|^{n+2s}}\big(-\ln|x-y|\big)\,dx\,dy$
$=c_{n,s}\iint_{\Omega\times\Omega}\frac{(u(x)-u(y))^{2}}{|x-y|^{n+2s}}\big(-\ln|x-y|\big)\,dx\,dy$
$\quad+2c_{n,s}\int_{\Omega}u(x)^{2}\left(\int_{\Omega^{c}}\frac{-\ln|x-y|}{|x-y|^{n+2s}}\,dy\right)\,dx.$
Fix $x\in\Omega$ and write $D:=\operatorname{dima}\Omega<1$. Since $|x-z|\leq D$ for every $z\in\Omega$, we have the inclusion $\Omega\subset B_{D}(x)$, hence
$\Omega^{c}\supset\mathbb{R}^{n}\setminus B_{D}(x).$
We split the exterior integral according to the sign of $\ln|x-y|$:
$\int_{\Omega^{c}}\frac{-\ln|x-y|}{|x-y|^{n+2s}}\,dy$ $=\int_{\Omega^{c}\cap\{|x-y|<1\}}\frac{-\ln|x-y|}{|x-y|^{n+2s}}\,dy-\int_{\Omega^{c}\cap\{|x-y|>1\}}\frac{\ln|x-y|}{|x-y|^{n+2s}}\,dy.$
Note that $\Omega^{c}\cap\{|x-y|<1\}\supset\{D<|x-y|<1\}$, we obtain
$\int_{\Omega^{c}\cap\{|x-y|<1\}}\frac{-\ln|x-y|}{|x-y|^{n+2s}}\,dy\ \geq\ \int_{D<|x-y|<1}\frac{-\ln|x-y|}{|x-y|^{n+2s}}\,dy=|\mathbb{S}^{n-1}|\int_{D}^{1}\frac{-\ln r}{r^{2s+1}}\,dr.$
Moreover, since $\Omega^{c}\cap\{|x-y|>1\}\subset\{|x-y|>1\}$ and the integrand is nonnegative,
$\int_{\Omega^{c}\cap\{|x-y|>1\}}\frac{\ln|x-y|}{|x-y|^{n+2s}}\,dy\ \leq\ \int_{|x-y|>1}\frac{\ln|x-y|}{|x-y|^{n+2s}}\,dy=|\mathbb{S}^{n-1}|\int_{1}^{\infty}\frac{\ln r}{r^{2s+1}}\,dr.$
Moreover, since $D<1,$
$\iint_{\Omega\times\Omega}\frac{(u(x)-u(y))^{2}}{|x-y|^{n+2s}}\big(-\ln|x-y|\big)\,dx\,dy\geq 0.$
Thus, we infer
$\mathcal{E}_{+}(u,u)-\mathcal{E}_{-}(u,u)\geq$ $\ 2c_{n,s}|\mathbb{S}^{n-1}|\Bigg(\int_{D}^{1}\frac{-\ln r}{r^{2s+1}}\,dr-\int_{1}^{\infty}\frac{\ln r}{r^{2s+1}}\,dr\Bigg)\|u\|_{L^{2}(\Omega)}^{2}$
$=$ $\frac{1}{s}c_{n,s}|\mathbb{S}^{n-1}|D^{-2s}(\ln\frac{1}{D}-\frac{1}{2s})\|u\|_{L^{2}(\Omega)}^{2}.$
(iii) For $\operatorname{dima}\Omega<e^{-\frac{1}{2s}}$ and $b_{n,s}\geq 0$, it is obvious that ${\mathcal{E}}_{s+\operatorname*{Log}}(u,u)\geq 0$ and ${\mathcal{E}}_{s+\operatorname*{Log}}(u,u)=0$ implies $u=0.$ By the proof of (ii), we obtain
${\mathcal{E}}_{+}(|u|,|u|)-{\mathcal{E}}_{-}(|u|,|u|)\leq{\mathcal{E}}_{+}(u,u)-{\mathcal{E}}_{-}(u,u).$
Therefore,
$0\leq{\mathcal{E}}_{s+\operatorname*{Log}}(|u|,|u|)$ $=\mathcal{E}_{+}(|u|,|u|)-\mathcal{E}_{-}(|u|,|u|)+\frac{b_{n,s}c_{n,s}}{2}{\mathcal{E}}_{s}(|u|,|u|)$
$\leq{\mathcal{E}}_{+}(u,u)-{\mathcal{E}}_{-}(u,u)+\frac{b_{n,s}c_{n,s}}{2}{\mathcal{E}}_{s}(u,u).$
Moreover, equality holds in (1.17) if and only if $u$ does not change sign. ∎
Next, we quantify the relation between the fractional-logarithmic energy and the pure fractional energy, making explicit the effect of the logarithmic factor; this estimate will be a key ingredient in proving compactness even at the critical exponent.
Lemma 3.1.
Let $0<r<1$. Then for every $u\in\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)$,
(3.4) $\mathcal{E}_{s}(u,u)\leq\frac{1}{c_{n,s}\,\ln(1/r)}\,\|u\|_{{\mathcal{H}}^{s+\operatorname*{Log}}_{0}(\Omega)}^{2}+\frac{2}{s}|\mathbb{S}^{n-1}|\,r^{-2s}\,\|u\|_{L^{2}(\Omega)}^{2}$
Proof.
Split the fractional energy into near and far fields:
$\mathcal{E}_{s}(u,u)=\iint_{|x-y|<r}\frac{(u(x)-u(y))^{2}}{|x-y|^{n+2s}}\,dx\,dy+\iint_{|x-y|\geq r}\frac{(u(x)-u(y))^{2}}{|x-y|^{n+2s}}\,dx\,dy=:I_{\rm near}+I_{\rm far}.$
For $|z|<r$, we have $(-\ln|z|)_{+}=\ln(1/|z|)\geq\ln(1/r)$, hence
$|z|^{-n-2s}\leq\frac{1}{\ln(1/r)}\,|z|^{-n-2s}(-\ln|z|)_{+}=\frac{1}{c_{n,s}\,\ln(1/r)}\,{\bf k}_{s+\operatorname*{Log},+}(z).$
Therefore,
$I_{\rm near}\leq\frac{1}{c_{n,s}\,\ln(1/r)}\iint_{\mathbb{R}^{n}\times\mathbb{R}^{n}}(u(x)-u(y))^{2}\,{\bf k}_{s+\operatorname*{Log},+}(x-y)\,dx\,dy=\frac{1}{c_{n,s}\,\ln(1/r)}\,[u]_{s+\operatorname*{Log},+}^{2}.$
For the far field, we use the integrability of $|z|^{-n-2s}$ on $\{|z|\geq r\}$:
$I_{\rm far}\leq$ $2\iint_{|x-y|\geq r}\frac{u(x)^{2}+u(y)^{2}}{|x-y|^{n+2s}}\,dx\,dy$
$=$ $4\int_{\mathbb{R}^{n}}u(x)^{2}\Bigl(\int_{|z|\geq r}|z|^{-n-2s}\,dz\Bigr)\,dx=\frac{2}{s}|\mathbb{S}^{n-1}|\,r^{-2s}\,\|u\|_{L^{2}(\Omega)}^{2}.$
Combining the two estimates yields (3.4). ∎
We now prove the embedding properties, precisely describing the relation between the fractional and fractional-logarithmic spaces, and showing that the embedding is compact for all $1\leq p\leq 2_{s}^{*}$.
Proof of Proposition 1.6.
(i) For any $\epsilon\in(0,1-s)$, there exists $r_{\epsilon}\in(0,e^{-\frac{1}{2s}})$ such that
$|z|^{-2s-n}(-2\ln|z|)_{+}\leq|z|^{-2(s+\epsilon)-n},\ 2s\ln r_{\epsilon}+1>0\quad{\rm for}\ \,0<|z|<r_{\epsilon}.$
Then a direct computation shows that
$\int_{\{|x-y|<r_{\epsilon}\}}\frac{\big(u(x)-u(y)\big)^{2}}{|x-y|^{n+2s}}(-2\ln|x-y|\big)_{+}dxdy$
$\leq$ $\int_{\{|x-y|<r_{\epsilon}\}}\frac{\big(u(x)-u(y)\big)^{2}}{|x-y|^{n+2s+2\epsilon}}dxdy\leq\|u\|_{{\mathcal{H}}^{s+\epsilon}_{0}(\Omega)}^{2}$
and
$\int_{\{|x-y|>r_{\epsilon}\}}\frac{\big(u(x)-u(y)\big)^{2}}{|x-y|^{n+2s}}(-\ln|x-y|\big)_{+}dxdy$
$\leq$ $4\int_{\mathbb{R}^{n}}u(x)^{2}\int_{\{y:|x-y|>r_{\epsilon}\}}\frac{(-\ln|x-y|\big)_{+}}{|x-y|^{n+2s}}dydx$
$=$ $4\left|\mathbb{S}^{n-1}\right|\left(\frac{1-r_{\epsilon}^{-2s}}{4s^{2}}\;-\;\frac{r_{\epsilon}^{-2s}\ln r_{\epsilon}}{2s}\right)\|u\|_{L^{2}(\Omega)}^{2}$
$\leq$ $c_{\epsilon}\|u\|_{L^{2}(\Omega)}^{2},$
where $c_{\epsilon}>0$ depends on $\epsilon$. Therefore, there exists $C_{\epsilon}>0$ such that
$\|u\|_{{\mathcal{H}}^{s+\operatorname*{Log}}_{0}(\Omega)}^{2}$ $=c_{n,s}\int_{\{|x-y|<r_{\epsilon}\}}\frac{\big(u(x)-u(y)\big)^{2}}{|x-y|^{n+2s}}(-\ln|x-y|\big)_{+}dxdy$
$+c_{n,s}\int_{\{|x-y|>r_{\epsilon}\}}\frac{\big(u(x)-u(y)\big)^{2}}{|x-y|^{n+2s}}(-\ln|x-y|\big)_{+}dxdy$
$\leq c_{n,s}\|u\|_{{\mathcal{H}}^{s+\epsilon}_{0}(\Omega)}^{2}+c_{\epsilon}c_{n,s}\|u\|_{L^{2}(\Omega)}^{2}$
$\leq C_{\epsilon}\|u\|_{{\mathcal{H}}^{s+\epsilon}_{0}(\Omega)}^{2}.$
That is, for any $\epsilon\in(0,1-s)$
${\mathcal{H}}^{s+\epsilon}_{0}(\Omega)\subset\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega).$
Note that
$\int_{\big\{|x-y|\geq\frac{1}{e}\big\}}\frac{(u(x)-u(y))^{2}}{|x-y|^{n+2s}}dxdy$
$\leq$ $4\int_{\mathbb{R}^{n}}u(x)^{2}\int_{\mathbb{R}^{n}}\frac{\chi_{(\frac{1}{e},+\infty)}(|x-y|)}{|x-y|^{n+2s}}dydx=\frac{2e^{2s}}{s}\left|\mathbb{S}^{n-1}\right|\|u\|_{L^{2}(\Omega)}^{2},$
where $\chi_{A}(t)=1$ if $t\in A$ and $\chi_{A}(t)=0$ if $t\not\in A$. Since $|z|^{-2s-n}(-\ln|z|)_{+}\geq|z|^{-2s-n}$ for $0<|z|<\frac{1}{e}$, then
$c_{n,s}\|u\|_{{\mathcal{H}}^{s}_{0}(\Omega)}^{2}$ $=c_{n,s}\iint_{\mathbb{R}^{n}\times\mathbb{R}^{n}}\frac{(u(x)-u(y))^{2}}{|x-y|^{n+2s}}\,dx\,dy$
$\leq\iint_{\mathbb{R}^{n}\times\mathbb{R}^{n}}(u(x)-u(y))^{2}\,{\bf k}_{s+\operatorname*{Log},+}(x-y)\,dx\,dy+$
$\qquad c_{n,s}\iint_{\{|x-y|\geq 1/e\}}\frac{(u(x)-u(y))^{2}}{|x-y|^{n+2s}}\,dx\,dy$
$\leq\iint_{\mathbb{R}^{n}\times\mathbb{R}^{n}}(u(x)-u(y))^{2}\,{\bf k}_{s+\operatorname*{Log},+}(x-y)\,dx\,dy+c_{n,s}\frac{2e^{2s}}{s}\,|\mathbb{S}^{n-1}|\,\|u\|_{L^{2}(\Omega)}^{2}.$
By Proposition 1.4, we obtain ${\mathcal{H}}^{s+\operatorname*{Log}}_{0}(\Omega)\ \hookrightarrow\ {\mathcal{H}}^{s}_{0}(\Omega).$
(ii) By the embedding properties of ${\mathcal{H}}^{s}_{0}(\Omega)$, it suffices to show that the embedding
${\mathcal{H}}^{s+\operatorname*{Log}}_{0}(\Omega)\hookrightarrow L^{2_{s}^{*}}(\Omega),$
is compact.
Fix $x_{0}\in\Omega,\ 0<r<\min\left\{\frac{1}{2}{\rm dist}(x_{0},\partial\Omega),1\right\}$ and $u\in{\mathcal{H}}^{s+\operatorname*{Log}}_{0}(\Omega).$ Choose $\eta\in C_{c}^{\infty}(B_{2r}(x_{0}))$ with $\eta\equiv 1$ on $B_{r}(x_{0})$ and $|\eta(x)-\eta(y)|\leq C\min\left\{1,|x-y|/r\right\},\left|\eta\right|\leq 1$. Set $v:=\eta u$, similar to the proof of Lemma 3.1, we have
$\mathcal{E}_{s}(v,v)\ \lesssim\ \frac{1}{\ln(1/r)}\mathcal{E}_{+}(u;B_{3r}(x_{0}))+r^{-2s}\,\|u\|_{L^{2}(\Omega)}^{2},$
where
$\mathcal{E}_{+}(u;B_{3r}(x_{0})):=\iint_{B_{3r}(x_{0})\times B_{3r}(x_{0})}(u(x)-u(y))^{2}\,{\bf k}_{s+\operatorname*{Log},+}(x-y)\,dx\,dy.$
By the fractional Sobolev inequality on $\mathbb{R}^{n}$ for $0<s<1$, $\|v\|_{L^{2_{s}^{*}}(\mathbb{R}^{n})}\lesssim\mathcal{E}_{s}(v,v)$, applied to $v=\eta u$ and using $\eta\equiv 1$ on $B_{r}(x_{0})$, we obtain
(3.5) $\|u\|_{L^{2_{s}^{*}}(B_{r}(x_{0}))}^{2}\ \lesssim\ \mathcal{E}_{s}(v,v)\ \lesssim\ \frac{1}{\ln(1/r)}\mathcal{E}_{+}(u;B_{3r}(x_{0}))+r^{-2s}\,\|u\|_{L^{2}(\Omega)}^{2}.$
Let $(u_{k})$ be a bounded sequence in ${\mathcal{H}}^{s+\operatorname*{Log}}_{0}(\Omega)$. By Proposition 1.6, $(u_{k})$ is bounded in $H^{s}_{0}(\Omega)$; the fractional Rellich theorem yields (up to a subsequence) $u_{k}\to u$ in $L^{2}(\Omega)$ and a.e. in $\Omega$ where $u\in{\mathcal{H}}^{s+\operatorname*{Log}}_{0}(\Omega).$ Fix $\varepsilon\in(0,1),$ choose $r$ so small that
(3.6) $\frac{1}{\ln(1/r)}\leq\varepsilon.$
Cover $\Omega$ by finitely many balls $\{B_{r}(x_{j})\}_{j=1}^{l}$ with bounded overlap.
Applying (3.5) to $u_{k}-u$ and summing over $j$, we get
$\|u_{k}-u\|_{L^{2_{s}^{*}}(\Omega)}^{2}\ \lesssim\ \frac{1}{\ln(1/r)}\mathcal{E}_{+}(u_{k}-u;\Omega)+r^{-2s}\,\|u_{k}-u\|_{L^{2}(\Omega)}^{2}.$
By the boundedness of $(u_{k})$ in ${\mathcal{H}}^{s+\operatorname*{Log}}_{0}(\Omega)$ and (3.6), the first term is $\lesssim\varepsilon$ uniformly in $k$; the second term tends to 0 as $k\to\infty$ because $u_{k}\to u$ in $L^{2}(\Omega)$. Hence
$\limsup_{k\to\infty}\|u_{k}-u\|_{L^{2_{s}^{*}}(\Omega)}^{2}\ \lesssim\ \varepsilon.$
Since $\varepsilon>0$ is arbitrary, $u_{k}\to u$ in $L^{2_{s}^{*}}(\Omega)$, i.e. the embedding ${\mathcal{H}}^{s+\operatorname*{Log}}_{0}(\Omega)\hookrightarrow L^{2_{s}^{*}}(\Omega)$ is compact. ∎
3.2. Existence and Regularity of Solutions
Next, we study the existence, uniqueness, and regularity properties of weak solutions to the nonlocal Poisson problem (1.18), governed by the fractional–logarithmic operator $(-\Delta)^{s+\operatorname*{Log}}$, where $w\in L^{\infty}(\Omega)$.
Definition 3.1.
Assume that $f\in\big(\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)\big)^{\!*}$ and $w\in L^{\infty}(\Omega)$. We say that $u\in\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)$ is a weak solution of (1.18) if
(3.7) $\mathcal{E}_{s+\operatorname*{Log}}(u,\varphi)+\int_{\Omega}V(x)\,u(x)\,\varphi(x)\,dx=\langle f,\varphi\rangle\qquad\text{for every }\varphi\in\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega),$
where $\langle\cdot,\cdot\rangle$ denotes the duality pairing between $\big(\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)\big)^{\!*}$ and $\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)$, and $\mathcal{E}_{s+\operatorname*{Log}}$ is defined in (1.15).
Lemma 3.2.
For every $\alpha>1,r>0$ and $u\in\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)$, there exists $C>0$ independent of $\alpha$ such that
[…]
Proof.
Note that
$\int_{\Omega}\int_{\mathbb{R}^{n}\setminus B_{r}(x)}|u|^{2\alpha-1}(x)\left|u(x)-u(y)\right|\left|{\bf K}_{s+\operatorname*{Log}}(|x-y|)\right|dydx$
$\leq$ $\int_{\Omega}\int_{\mathbb{R}^{n}\setminus B_{r}(x)}|u|^{2\alpha-1}(x)\left(\left|u(x)\right|+\left|u(y)\right|\right)\left|{\bf K}_{s+\operatorname*{Log}}(|x-y|)\right|dydx$
$=$ $\int_{\Omega}\int_{\mathbb{R}^{n}\setminus B_{r}(x)}u^{2\alpha}(x){\bf K}_{s+\operatorname*{Log}}(|x-y|)dydx+\int_{\Omega}\int_{\mathbb{R}^{n}\setminus B_{r}(x)}|u^{2\alpha-1}(x)||u(y)|{\bf K}_{s+\operatorname*{Log}}(|x-y|)dydx$
[…]
By the property of ${\bf K}_{s+\operatorname*{Log}}$, we can obtain that
$\int_{r}^{\infty}\rho^{n-1}{\bf K}_{s+\operatorname*{Log}}(\rho)d\rho<\infty,\quad\int_{\mathbb{R}^{n}\setminus B_{r}(x)}\left|{\bf K}_{s+\operatorname*{Log}}(|x-y|)\right|^{2}dy<\infty.$
combining the fact
[…]
thus, we obtain that (3.8) holds. ∎
We now use the Lax–Milgram theorem to prove existence (and uniqueness) of weak solutions, and then apply a Moser iteration to establish $L^{\infty}$-regularity.
Proof of Theorem 1.4.
(i) Define the bilinear form
$a(u,\varphi):={\mathcal{E}}_{s+\operatorname*{Log}}(u,\varphi)+\int_{\Omega}V(x)\,u(x)\,\varphi(x)\,dx,\quad u,\varphi\in\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega),$
and the linear functional $\ell(\varphi):=\langle f,\varphi\rangle$. By Cauchy–Schwarz inequality, the boundedness of $w$, and the estimates in Proposition 1.6 (i), $a$ is continuous on $\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)$ and $\ell$ is continuous on $\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)$.
Next we prove coercivity. For $u\in\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)$ we write
$a(u,u)$ $=[u]_{s+\operatorname*{Log},+}^{2}-{\mathcal{E}}_{-}(u,u)+\frac{b_{n,s}c_{n,s}}{2}\,{\mathcal{E}}_{s}(u,u)+\int_{\Omega}V\,u^{2}\,dx$
$\geq[u]_{s+\operatorname*{Log},+}^{2}-{\mathcal{E}}_{-}(u,u)-\frac{|b_{n,s}|c_{n,s}}{2}\,{\mathcal{E}}_{s}(u,u)+\big(\operatorname{essinf}_{\Omega}V\big)\|u\|_{L^{2}(\Omega)}^{2}.$
Using Proposition 1.5 (i) we have
${\mathcal{E}}_{-}(u,u)\leq\frac{c_{n,s}|\mathbb{S}^{n-1}|}{s^{2}}\,\|u\|_{L^{2}(\Omega)}^{2},$
and by Lemma 3.1 together with $\|u\|_{\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)}=[u]_{s+\operatorname*{Log},+}$ we obtain
${\mathcal{E}}_{s}(u,u)\leq\frac{1}{c_{n,s}\ln(1/r)}\,\|u\|_{\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)}^{2}+\frac{2}{s}|\mathbb{S}^{n-1}|\,r^{-2s}\,\|u\|_{L^{2}(\Omega)}^{2}.$
Plugging these estimates into the previous inequality yields
$a(u,u)$ $\geq\Bigl(1-\frac{|b_{n,s}|}{2\,\ln(1/r)}\Bigr)\|u\|_{\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)}^{2}+$
$\quad\Bigl(\operatorname{essinf}_{\Omega}V-\frac{c_{n,s}|\mathbb{S}^{n-1}|}{s^{2}}-\frac{|b_{n,s}|c_{n,s}}{s}\,|\mathbb{S}^{n-1}|\,r^{-2s}\Bigr)\|u\|_{L^{2}(\Omega)}^{2}$
$=\alpha_{r}\,\|u\|_{\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)}^{2}+\Bigl(\operatorname{essinf}_{\Omega}V-\frac{c_{n,s}|\mathbb{S}^{n-1}|}{s^{2}}-\frac{|b_{n,s}|c_{n,s}}{s}\,|\mathbb{S}^{n-1}|\,r^{-2s}\Bigr)\|u\|_{L^{2}(\Omega)}^{2}.$
By assumption (1.20),
$a(u,u)\geq\alpha_{r}\|u\|_{\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)}^{2}.$
Therefore, the Lax–Milgram theorem yields a unique $u\in\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)$ such that $a(u,\varphi)=\ell(\varphi)$ for all $\varphi\in\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)$, which is exactly the weak formulation of (1.18). The priori bound follows from the standard Lax–Milgram estimate.
(ii) Without loss of generality, we only consider $u$ is non-negative. For $\alpha>1$ and $R>0$ large, we take
$\varphi_{R}(x)=\begin{cases}x^{\alpha},&\text{if }0\leq x<R,\\[6.0pt] \alpha R^{\alpha-1}x-(\alpha-1)R^{\alpha},&\text{if }x\geq R.\end{cases}$
It is esay to see that $\varphi_{R}$ is convex and Lipschitz function with $\varphi_{R}(0)=0.$ Thus $\varphi_{R}\left(u\right)\in\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega)$ since $u\in\mathcal{H}^{s+\operatorname*{Log}}_{0}(\Omega).$ By Proposition 1.6, there exists $c_{1},c_{2}>0$ satisfying
[…]
$\leq c_{2}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\left|\varphi_{R}\left(u\right)(x)-\varphi_{R}\left(u\right)(y)\right|^{2}\left|{\bf K}_{s+\operatorname*{Log}}(x-y)\right|dxdy$
$=2\int_{\Omega}\varphi_{R}\left(u\right)(x)\left|\left(-\Delta\right)^{s+\operatorname*{Log}}\right|\varphi_{R}\left(u\right)(x)dx$
where
$\left|\left(-\Delta\right)^{s+\operatorname*{Log}}\right|u(x):=\int_{\mathbb{R}^{n}}\left(u(x)-u(y)\right)\left|{\bf K}_{s+\operatorname*{Log}}(x-y)\right|dy.$
Note that there exists a constant $\eta>0$ such that
$|{\bf K}_{s+\operatorname*{Log}}(r)|={\bf K}_{s+\operatorname*{Log}}(r)\quad\text{for all }r\in(0,\eta),$
so by the convexity of $\varphi_{R}$,
$\left|\left(-\Delta\right)^{s+\operatorname*{Log}}\right|\varphi_{R}\left(u\right)(x)\leq\left|\varphi_{R}^{\prime}(u)(x)\right|\left|\left(-\Delta\right)^{s+\operatorname*{Log}}\right|u(x)$
$\leq$ $\left|\varphi_{R}^{\prime}(u)(x)\right|\left(-\Delta\right)^{s+\operatorname*{Log}}u(x)+2\left|\varphi_{R}^{\prime}(u)(x)\right|\int_{B_{\eta}\left(x\right)^{c}}\left|u(x)-u(y)\right|\left|{\bf K}_{s+\operatorname*{Log}}(|x-y|)\right|dy$
By Lemma 3.2 and the fact $\varphi_{R}(x)\leq x^{\alpha},\varphi_{R}^{\prime}\left(x\right)\leq\alpha x^{\alpha-1},$ there exists $c_{3}>0$ such that
[…]
thus, there exists $c_{4}>0$ such that
[…]
Take $p=\frac{2n}{n-2s}$ and $k=\frac{2q}{q-1}$, then $p>k>2$. Let $R\rightarrow\infty$, we obtain that
[…]
where $c_{2}>\max\left\{1,|\Omega\right\}.$ Note that
[…]
thus,
$\left(\int_{\Omega}u^{\alpha p}(x)dx\right)^{\frac{1}{\alpha p}}\leq C_{\alpha}^{\frac{1}{2\alpha}}\left(c_{2}+\int_{\Omega}u^{k\alpha}dx\right)^{\frac{1}{k\alpha}},$
where
[…]
Next, we proceed by employing a standard iteration argument. Define
$\alpha_{0}=1,\alpha_{i}=\left(\frac{p}{k}\right)^{i}>1,$
then $\alpha_{i+1}k=\alpha_{i}p$. Denote
$A_{i}=\left(\int_{\Omega}u^{\alpha_{i}p}(x)dx\right)^{\frac{1}{\alpha_{i}p}},\ C_{i}=C_{\alpha_{i}}^{\frac{1}{2\alpha_{i}}}.$
Thus, we obtain that
$A_{i+1}\leq C_{i+1}\left(c_{2}+A_{i}^{\alpha_{i}p}\right)^{\frac{1}{\alpha_{i}p}}.$
Without loss of generality, we suppose $A_{0}=1$ and $|\Omega|=1,$ then $A_{i}\geq 1,\forall\ i\geq 1.$ Taking logarithms and
$\log A_{i+1}\leq\log C_{i+1}+\frac{1}{\alpha_{i}p}\log\left(c_{2}+A_{i}^{\alpha_{i}p}\right)\leq\log C_{i+1}+\frac{c_{2}}{\alpha_{i}p}+\log A_{i}.$
Thus, we have
$\log A_{i+1}\leq\sum_{j=1}^{i+1}\log C_{j}+\sum_{j=0}^{i}\frac{c_{2}}{\alpha_{j}p},$
by the fact $p>k>2$ and
$\sum_{j=1}^{i+1}\log C_{j}=\sum_{j=1}^{i+1}\frac{1}{2\alpha_{j}}\log C_{\alpha_{j}}<\infty,$
thus, there exists a constant $K>0$ such that
$\|u\|_{L^{q}}\leq K,\quad 1\leq q<\infty.$
Therefore, we obtain $u\in L^{\infty}(\Omega).$
4. The Dirichlet eigenvalue problem
In this section, we investigate the Dirichlet eigenvalue problem associated with the fractional-logarithmic operator. We establish the variational framework and derive the min–max characterization of eigenvalues, together with basic spectral properties such as orthogonality and completeness of eigenfunctions, and divergence of the eigenvalue sequence. We then turn to high-frequency behavior and prove a Weyl-type asymptotic law for the counting function and for the $k$-th eigenvalue.
We first establish the basic properties of the Dirichlet eigenvalues.
Proof of Theorem 1.5..
By Proposition 1.5 and Proposition 1.6, one easily verifies that the functional $\Phi_{s+\operatorname*{Log}}:{\mathcal{H}}^{s+\operatorname*{Log}}_{0}(\Omega)\to\mathbb{R}$ defined by
$\Phi_{s+\operatorname*{Log}}(u):={\mathcal{E}}_{s+\operatorname*{Log}}(u,u)\qquad\text{for every $u\in{\mathcal{H}}^{s+\operatorname*{Log}}_{0}(\Omega)$}$
is weakly lower semicontinuous (in fact, it is even of the class $C^{1}$). Thus, the remaining proof of Theorem 1.5 is similar to the one of [7, Theorem 1.4]. For the convenience of the readers, we provide here the details.
(i) Let
${\mathcal{M}}_{1}:=\{u\in{\mathcal{H}}^{s+\operatorname*{Log}}_{0}(\Omega),\,\|u\|_{L^{2}(\Omega)}=1\}.$
Then, we note that
(4.1) ${\mathcal{E}}_{s+\operatorname*{Log}}(u,u)\geq\|u\|_{{\mathcal{H}}^{s+\operatorname*{Log}}_{0}(\Omega)}-C_{s}\left(1+\|u\|_{{\mathcal{H}}^{s+\operatorname*{Log}}_{0}(\Omega)}\right)>$-$\infty$
for every $u\in{\mathcal{M}}_{1}$. Thus, and by the lower semicontinuity of $\Phi_{s+\operatorname*{Log}}$, the first eigenvalue
$\lambda_{1}^{s+\operatorname*{Log}}(\Omega):=\inf_{{\mathcal{M}}_{1}}\Phi_{s+\operatorname*{Log}}$
is attained by a function $\xi_{1}\in{\mathcal{M}}_{1}$. Consequently, there exists a Lagrange multiplier $\lambda\in\mathbb{R}$ such that
${\mathcal{E}}_{s+\operatorname*{Log}}(\xi_{1},\varphi)=\lambda\int_{\Omega}\xi_{1}\varphi\,dx\qquad\text{for all $\varphi\in{\mathcal{H}}^{s+\operatorname*{Log}}_{0}(\Omega)$.}$
Choosing $\varphi=\xi_{1}$ yields $\lambda=\lambda_{1}^{s+\operatorname*{Log}}(\Omega)$, hence $\xi_{1}$ is an eigenfunction of (1.21) corresponding to $\lambda_{1}^{s+\operatorname*{Log}}(\Omega)$.
Next we proceed inductively and assume that $\xi_{2},\dots,\xi_{k}\in{\mathcal{H}}^{s+\operatorname*{Log}}_{0}(\Omega)$ and
$\lambda_{2}^{s+\operatorname*{Log}}(\Omega)\leq\dots\leq\lambda_{k}^{s+\operatorname*{Log}}(\Omega)$
are already given for some $k\in\mathbb{N}$ with the properties that for $i=2,\dots,k$, the function $\xi_{i}$ is a minimizer of ${\mathcal{E}}_{s+\operatorname*{Log}}(\cdot,\cdot)$ within the set
${\mathcal{M}}_{i}$ $:=\big\{u\in{\mathcal{H}}_{i}(\Omega)\>:\>\|u\|_{L^{2}(\Omega)}=1\big\}$
$=\Big\{u\in{\mathcal{H}}^{s+\operatorname*{Log}}_{0}(\Omega)\>:\>\|u\|_{L^{2}(\Omega)}=1,\>\text{$\int_{\Omega}u\xi_{j}\,dx=0$ for $j=1,\dots i-1$}\Big\}.$
Moreover,
$\lambda_{i}^{s+\operatorname*{Log}}(\Omega)=\inf_{u\in{\mathcal{M}}_{i}}\Phi_{s+\operatorname*{Log}}(u)=\Phi_{s+\operatorname*{Log}}(\xi_{i}),$
and
(4.2) ${\mathcal{E}}_{s+\operatorname*{Log}}(\xi_{i},\varphi)=\lambda_{i}^{s+\operatorname*{Log}}(\Omega)\int_{\Omega}\xi_{i}\varphi\,dx\qquad\text{for all $\varphi\in{\mathcal{H}}^{s+\operatorname*{Log}}_{0}(\Omega)$.}$
We then set
${\mathcal{H}}_{k+1}(\Omega):=\left\{u\in{\mathcal{H}}^{s+\operatorname*{Log}}_{0}(\Omega)\>:\>\|u\|_{L^{2}(\Omega)}=1,\>\text{$\int_{\Omega}u\xi_{i}\,dx=0$ for $i=1,\dots k$}\right\},$
and
${\mathcal{M}}_{k+1}:=\big\{u\in{\mathcal{H}}_{k+1}(\Omega),\,\|u\|_{L^{2}(\Omega)}=1\big\},\quad\lambda_{k+1}^{s+\operatorname*{Log}}(\Omega):=\inf_{u\in{\mathcal{M}}_{k+1}}\Phi_{s+\operatorname*{Log}}(u).$
By the same weak lower semicontinuity argument as above, the value $\lambda_{k+1}^{s+\operatorname*{Log}}(\Omega)$ is attained by a function $\xi_{k+1}\in{\mathcal{M}}_{k+1}$. Consequently, there exists a Lagrange multiplier $\lambda\in\mathbb{R}$ with the property that
(4.3) ${\mathcal{E}}_{s+\operatorname*{Log}}(\xi_{k+1},\varphi)=\lambda\int_{\Omega}\xi_{k+1}\varphi\,dx\qquad\text{for all $\varphi\in{\mathcal{H}}_{k+1}(\Omega)$.}$
Choosing $\varphi=\xi_{k+1}$, it yields $\lambda=\lambda_{k+1}^{s+\operatorname*{Log}}(\Omega)$. Moreover, for $i=1,\dots,k$, we have, by (4.2) and the definition of ${\mathcal{H}}_{k+1}(\Omega)$,
${\mathcal{E}}_{s+\operatorname*{Log}}(\xi_{k+1},\xi_{i})$ $={\mathcal{E}}_{s+\operatorname*{Log}}(\xi_{i},\xi_{k+1})$
$=\lambda_{i}^{s+\operatorname*{Log}}(\Omega)\int_{\Omega}\xi_{i}\xi_{k+1}\,dx=0=\lambda_{k+1}^{s+\operatorname*{Log}}(\Omega)\int_{\Omega}\xi_{k+1}\xi_{i}\,dx.$
Hence (4.3) holds with $\lambda=\lambda_{k+1}^{s+\operatorname*{Log}}(\Omega)$ for all $\varphi\in{\mathcal{H}}^{s+\operatorname*{Log}}_{0}(\Omega)$.
Therefore, we have constructed an $L^{2}$-normalized sequence $(\xi_{k})_{k}$ in ${\mathcal{H}}^{s+\operatorname*{Log}}_{0}(\Omega)$ and a nondecreasing sequence $(\lambda_{k}^{s+\operatorname*{Log}}(\Omega))_{k}$ in $\mathbb{R}$ such that property (i) holds and such that $\xi_{k}$ is an eigenfunction of (1.21) corresponding to $\lambda=\lambda_{k}^{s+\operatorname*{Log}}(\Omega)$.
(ii) Supposing by contradiction that $c:=\lim\limits_{k\to\infty}\lambda_{k}^{s+\operatorname*{Log}}(\Omega)<+\infty$, we deduce that $\Phi_{s+\operatorname*{Log}}(\xi_{k})\leq c$ for every $k\in\mathbb{N}$. By Lemma 3.1, we can obtain that the sequence $(\xi_{k})$ is bounded in ${\mathcal{H}}^{s+\operatorname*{Log}}_{0}(\Omega)$, and therefore it contains a convergent subsequence $(\xi_{k_{j}})_{j}$ in $L^{2}(\Omega)$ by the compactness of embedding. This however is impossible since the functions $\{\xi_{k_{j}}\}_{j\in\mathbb{N}}$ are $L^{2}$-orthonormal. Hence (ii) is proved.
(iii) We first suppose by contradiction that there exists $v\in{\mathcal{H}}^{s+\operatorname*{Log}}_{0}(\Omega)$ with $\|v\|_{L^{2}(\Omega)}=1$ and $\int_{\Omega}v\xi_{k}dx=0$ for any $k\in\mathbb{N}$. Since $\lim\limits_{k\to\infty}\lambda_{k}^{s+\operatorname*{Log}}(\Omega)=+\infty$, there exists an integer $k_{0}>0$ such that
$\Phi_{s+\operatorname*{Log}}(v)<\lambda_{k_{0}}^{s+\operatorname*{Log}}(\Omega)=\inf_{{\mathcal{M}}_{k_{0}}}\Phi_{s+\operatorname*{Log}}(u),$
which by definition of ${\mathcal{M}}_{k_{0}}$ implies that $\int_{\Omega}v\xi_{k}dx\not=0$ for some $k\in\{1,\dots,k_{0}-1\}$. Therefore, we conclude that ${\mathcal{H}}^{s+\operatorname*{Log}}_{0}(\Omega)$ is contained in the $L^{2}$-closure of the span of $\{\xi_{k}\>:\>k\in\mathbb{N}\}$. Since ${\mathcal{H}}^{s+\operatorname*{Log}}_{0}(\Omega)$ is dense in $L^{2}(\Omega)$, we conclude that the span of $\{\xi_{k}\>:\>k\in\mathbb{N}\}$ is dense in $L^{2}(\Omega)$, and hence $\{\xi_{k}\>:\>k\in\mathbb{N}\}$ is an orthonormal basis.
(iv) Let $w\in{\mathcal{H}}^{s+\operatorname*{Log}}_{0}(\Omega)$ be a $L^{2}$-normalized eigenfunction of (1.21) corresponding to the eigenvalue $\lambda_{1}^{s+\operatorname*{Log}}(\Omega)$, i.e. we have
(4.4) ${\mathcal{E}}_{s+\operatorname*{Log}}(w,\varphi)=\lambda_{1}^{s+\operatorname*{Log}}(\Omega)\int_{\Omega}w\varphi\,dx\qquad\text{for all $w\in{\mathcal{H}}^{s+\operatorname*{Log}}_{0}(\Omega)$.}$
We show that $w$ does not change sign. Indeed, choosing $\varphi=w$ in (4.4), we see that $w$ is a minimizer of ${\mathcal{E}}_{s+\operatorname*{Log}}\big|_{{\mathcal{M}}_{1}}$. On the other hand, we also have $|w|\in{\mathcal{M}}_{1}$ and
${\mathcal{E}}_{s+\operatorname*{Log}}(|w|,|w|)\leq{\mathcal{E}}_{s+\operatorname*{Log}}(w,w)=\lambda_{1}^{s+\operatorname*{Log}}(\Omega).$
Hence equality holds by definition of $\lambda_{1}^{s+\operatorname*{Log}}(\Omega)$, and then Proposition 1.5 (iii) implies that $w$ does not change sign. Therefore, we may assume that $\xi_{1}$ is nonnegative, by Proposition 1.5, we obatin $\lambda_{1}^{s+\operatorname*{Log}}(\Omega)>0.$
Next, we introduce the Riesz mean of the shifted operator (see [16]):
${\rm Tr}\big((-\Delta)^{s+\operatorname*{Log}}-\Lambda\big)_{-}=\sum_{k\in\mathbb{N}}\big(\lambda_{k}^{s+\operatorname*{Log}}(\Omega)-\Lambda\big)_{-},\qquad\Lambda\in\mathbb{R}.$
We first derive an asymptotic upper bound for the Riesz mean of the Dirichlet spectrum.
Lemma 4.1.
Let $\Omega$ be a bounded domain. Then one has that
(4.5) ${\rm Tr}\big((-\Delta)^{s+\operatorname*{Log}}-\Lambda\big)_{-}\leq\frac{|\Omega|}{(2\pi)^{n}}\frac{2|\mathbb{S}^{n-1}|}{n(n+2s)}(s\Lambda)^{1+\frac{n}{2s}}\big(\ln s\Lambda\big)^{-\frac{n}{2s}}\big(1+o(1)\big)$
as $\Lambda\rightarrow\infty$.
Proof.
It follows by [14, Theorem 2] that
${\rm Tr}\big((-\Delta)^{s+\operatorname*{Log}}-\Lambda\big)_{-}\leq\frac{|\Omega|}{(2\pi)^{n}}\int_{\mathbb{R}^{n}}\big(|\xi|^{2s}\ln|\xi|^{2}-\Lambda\big)_{-}d\xi.$
By the theory of Lambert $W$ function, a direct computation shows that for given $\Lambda>0$,
$r_{\Lambda}^{2s}\ln r_{\Lambda}^{2}=\Lambda\qquad\Longrightarrow\qquad r_{\Lambda}\leq(s\Lambda)^{\frac{1}{2s}}\big(\ln\Lambda\big)^{-\frac{1}{2s}}\Big(1+\frac{1}{2s}\frac{\ln\ln\Lambda}{\ln\Lambda}\big(1+o(1)\big)\Big),$
which implies that for sufficiently large $\Lambda$
$\int_{\mathbb{R}^{n}}\big(|\xi|^{2s}\ln|\xi|^{2}-\Lambda\big)_{-}d\xi$ $=\int_{B_{r_{\Lambda}}}\big(\Lambda-|\xi|^{2s}\ln|\xi|^{2}\big)d\xi$
$=|\mathbb{S}^{n-1}|\Big(\frac{1}{n}\Lambda r_{\Lambda}^{n}-\frac{2}{n+2s}r_{\Lambda}^{2s+n}\big(\ln r_{\Lambda}-\frac{1}{n+2s}\big)\Big)$
$=|\mathbb{S}^{n-1}|\left(\frac{2s\Lambda r_{\Lambda}^{n}}{n\left(n+2s\right)}+\frac{2}{\left(n+2s\right)^{2}}r_{\Lambda}^{2s+n}\right)$
$=\frac{2|\mathbb{S}^{n-1}|}{n(n+2s)}(s\Lambda)^{1+\frac{n}{2s}}\big(\ln\Lambda\big)^{-\frac{n}{2s}}\big(1+o(1)\big).$
The proof ends. ∎
With this lemma at hand, we can now complete the proof.
Proof of Theorem 1.6..
Let $g\in C_{c}^{\infty}(\mathbb{R}^{n})$ be a real-valued, $L^{2}$-normalized function with support in $\{x\in\mathbb{R}^{n}:|x|\leq\frac{\delta}{2}\}$. For $p\in\mathbb{R}^{n}$ and $q\in\Omega_{\delta}$, the coherent states
$F_{p,q}(x)=e^{{\rm i}p\cdot x}g(x-q),$
where $\Omega_{\delta}=\{x\in\Omega:\,{\rm dist}(x,\mathbb{R}^{n}\setminus\Omega)>\delta\}$.
From [17, Theorem 12.8], there holds
${\rm Tr}((-\Delta)^{s+\operatorname*{Log}}-\Lambda)_{-}$ $\geq(2\pi)^{-n}\int\int_{\mathbb{R}^{n}\times\Omega_{\delta}}\big\langle F_{p,q},\big((-\Delta)^{s+\operatorname*{Log}}-\Lambda\big)_{-}F_{p,q}\big\rangle\,dpdq$
$\geq(2\pi)^{-n}\int\int_{\mathbb{R}^{n}\times\Omega_{\delta}}\Big(\big\langle F_{p,q},(-\Delta)^{s+\operatorname*{Log}}F_{p,q}\big\rangle-\Lambda\Big)_{-}\,dpdq.$
Similar to the proof of [14, (8)], there exists $C>0$ such that
$\langle F_{p,q},(-\Delta)^{s+\operatorname*{Log}}F_{p,q}\rangle\leq|p|^{2s}\ln|p|^{2}+C,$
thus
${\rm Tr}((-\Delta)^{s+\operatorname*{Log}}-\Lambda)_{-}\geq$ $\frac{|\Omega_{\delta}|}{(2\pi)^{n}}\int_{\mathbb{R}^{n}}\Big(|p|^{2s}\ln|p|^{2}+C-\Lambda\Big)_{-}dp$
$=$ $\frac{|\Omega_{\delta}|}{(2\pi)^{n}}\frac{2|\mathbb{S}^{n-1}|}{n(n+2s)}(s\Lambda)^{1+\frac{n}{2s}}\big(\ln\Lambda\big)^{-\frac{n}{2s}}\big(1+o(1)\big),$
which, together with (4.5), implies that
$\lim_{\Lambda\to+\infty}{\rm Tr}\big((-\Delta)^{s+\operatorname*{Log}}-\Lambda\big)_{-}\Big(\Lambda^{-1-\frac{n}{2s}}\big(\ln\Lambda\big)^{\frac{n}{2s}}\Big)=\frac{|\Omega|}{(2\pi)^{n}}\frac{2|\mathbb{S}^{n-1}|}{n(n+2s)}s^{1+\frac{n}{2s}}=:A_{s}.$
Let ${\mathcal{S}}_{0}(\Lambda)=\sum_{k}(\Lambda-\lambda_{k}^{s+\operatorname*{Log}}(\Omega))_{+}={\rm Tr}\big((-\Delta)^{s+\operatorname*{Log}}-\Lambda\big)_{-}$, then for any $\kappa\in(0,\frac{\Lambda}{2})$,
$\big(\Lambda+\kappa-\lambda_{k}^{s+\operatorname*{Log}}(\Omega)\big)_{+}-\big(\Lambda-\lambda_{k}^{s+\operatorname*{Log}}(\Omega)\big)_{+}\geq\kappa\big(\Lambda-\lambda_{k}^{s+\operatorname*{Log}}(\Omega)\big)_{+}^{0}=\kappa{\mathcal{N}}(\Lambda)$
and
$\big(\Lambda-\lambda_{k}^{s+\operatorname*{Log}}(\Omega)\big)_{+}-\big(\Lambda-\kappa-\lambda_{k}^{s+\operatorname*{Log}}(\Omega)\big)_{+}\leq\kappa\big(\Lambda-\lambda_{k}^{s+\operatorname*{Log}}(\Omega)\big)_{+}^{0}=\kappa{\mathcal{N}}(\Lambda),$
that is,
(4.6) ${\mathcal{S}}_{0}(\Lambda)-{\mathcal{S}}_{0}(\Lambda-\kappa)\leq\kappa{\mathcal{N}}(\Lambda)\leq{\mathcal{S}}_{0}(\Lambda+\kappa)-{\mathcal{S}}_{0}(\Lambda).$
Note that the function
$F(\Lambda):=\Lambda^{1+\frac{n}{2s}}(\ln\Lambda)^{-\frac{n}{2s}},\qquad\Lambda>1,$
is increasing and convex on $(\Lambda_{1},\infty)$ for some sufficiently large $\Lambda_{1}$. Fix $\epsilon\in(0,\tfrac{1}{2})$. By the asymptotic relation in (4.6), there exists $\Lambda_{2}>2\Lambda_{1}$ such that for all $\Lambda\geq\Lambda_{2}$,
$\Big|\big(\Lambda^{-1-\frac{n}{2s}}(\ln\Lambda)^{\frac{n}{2s}}\big){\mathcal{S}}_{0}(\Lambda)-A_{s}\Big|\leq\epsilon,$
equivalently,
$(A_{s}-\epsilon)\,F(\Lambda)\leq{\mathcal{S}}_{0}(\Lambda)\leq(A_{s}+\epsilon)\,F(\Lambda).$
Take $\kappa=\sqrt{\epsilon}\,\Lambda$. Then $\Lambda+\kappa=\Lambda(1+\sqrt{\epsilon})\geq\Lambda_{1}$, so that $F$ is increasing and convex on the interval $\Lambda,\Lambda+\kappa]$. Applying (4.6) and the above bounds yields
${\mathcal{N}}(\Lambda)\leq\frac{1}{\kappa}\Big\{A_{s}\big(F(\Lambda+\kappa)-F(\Lambda)\big)+\epsilon\big(F(\Lambda+\kappa)+F(\Lambda)\big)\Big\}.$
Since $F$ is convex, we have for any $\kappa>0$,
$\frac{F(\Lambda+\kappa)-F(\Lambda)}{\kappa}\leq F^{\prime}(\Lambda+\kappa),$
while the monotonicity of $F$ implies $F(\Lambda)\leq F(\Lambda+\kappa)$ and hence
$\frac{\epsilon}{\kappa}\big(F(\Lambda+\kappa)+F(\Lambda)\big)\leq\frac{2\epsilon}{\kappa}F(\Lambda+\kappa)=\frac{2\sqrt{\epsilon}}{\Lambda}\,F(\Lambda+\kappa).$
Moreover, a direct computation gives
$F^{\prime}(\Lambda)=\frac{n+2s}{2s}\,\Lambda^{\frac{n}{2s}}(\ln\Lambda)^{-\frac{n}{2s}-1}\Bigl(\ln\Lambda-\frac{n}{n+2s}\Bigr).$
Combining these estimates and using $\Lambda+\kappa=\Lambda(1+\sqrt{\epsilon})$, we obtain
${\mathcal{N}}(\Lambda)$ $\leq A_{s}\,F^{\prime}(\Lambda+\kappa)+\frac{2\sqrt{\epsilon}}{\Lambda}\,F(\Lambda+\kappa)$
$\leq\frac{n+2s}{2s}A_{s}\,[\Lambda(1+\sqrt{\epsilon})]^{\frac{n}{2s}}\bigl(\ln[\Lambda(1+\sqrt{\epsilon})]\bigr)^{-\frac{n}{2s}-1}\Bigl(\ln[\Lambda(1+\sqrt{\epsilon})]-\frac{n}{n+2s}\Bigr)$
$\quad+\frac{2\sqrt{\epsilon}}{\Lambda}\,\{\Lambda(1+\sqrt{\epsilon})\}^{1+\frac{n}{2s}}\bigl(\ln[\Lambda(1+\sqrt{\epsilon})]\bigr)^{-\frac{n}{2s}}.$
and similarly we derive a lower bound. With the same notation, we obtain
${\mathcal{N}}(\Lambda)\geq\frac{1}{\kappa}\Big\{A_{s}\big(F(\Lambda)-F(\Lambda-\kappa)\big)-\epsilon\big(F(\Lambda)+F(\Lambda-\kappa)\big)\Big\}.$
By convexity of $F$, the derivative is increasing, hence
$\frac{F(\Lambda)-F(\Lambda-\kappa)}{\kappa}\geq F^{\prime}(\Lambda-\kappa).$
Moreover, by monotonicity $F(\Lambda-\kappa)\leq F(\Lambda)$ and thus
$\frac{\epsilon}{\kappa}\big(F(\Lambda)+F(\Lambda-\kappa)\big)\leq\frac{2\epsilon}{\kappa}F(\Lambda)=2\sqrt{\epsilon}\,\Lambda^{\frac{n}{2s}}(\ln\Lambda)^{-\frac{n}{2s}}.$
Using the explicit expression
$F^{\prime}(\Lambda)=\frac{n+2s}{2s}\,\Lambda^{\frac{n}{2s}}(\ln\Lambda)^{-\frac{n}{2s}-1}\Bigl(\ln\Lambda-\frac{n}{n+2s}\Bigr),$
and $\Lambda-\kappa=\Lambda(1-\sqrt{\epsilon})$, we arrive at
${\mathcal{N}}(\Lambda)$ $\geq A_{s}\,F^{\prime}(\Lambda-\kappa)-2\sqrt{\epsilon}\,\Lambda^{\frac{n}{2s}}(\ln\Lambda)^{-\frac{n}{2s}}$
$\geq\frac{n+2s}{2s}A_{s}\,[\Lambda(1-\sqrt{\epsilon})]^{\frac{n}{2s}}\bigl(\ln[\Lambda(1-\sqrt{\epsilon})]\bigr)^{-\frac{n}{2s}-1}\Bigl(\ln[\Lambda(1-\sqrt{\epsilon})]-\frac{n}{n+2s}\Bigr)$
$\quad-2\sqrt{\epsilon}\,\Lambda^{\frac{n}{2s}}(\ln\Lambda)^{-\frac{n}{2s}}.$
Dividing the upper and lower bounds by $\Lambda^{\frac{n}{2s}}(\ln\Lambda)^{-\frac{n}{2s}}$ and letting $\Lambda\to\infty$, we obtain
$\limsup_{\Lambda\to\infty}\Bigl(\Lambda^{-\frac{n}{2s}}(\ln\Lambda)^{\frac{n}{2s}}{\mathcal{N}}(\Lambda)\Bigr)\leq\frac{n+2s}{2s}A_{s}\,(1+o_{\epsilon}(1))+C\sqrt{\epsilon},$
and
$\liminf_{\Lambda\to\infty}\Bigl(\Lambda^{-\frac{n}{2s}}(\ln\Lambda)^{\frac{n}{2s}}{\mathcal{N}}(\Lambda)\Bigr)\geq\frac{n+2s}{2s}A_{s}\,(1+o_{\epsilon}(1))-C\sqrt{\epsilon},$
where $o_{\epsilon}(1)\to 0$ as $\Lambda\to\infty$ for each fixed $\epsilon$, and $C$ is independent of $\Lambda$ and $\epsilon$. Since $\epsilon\in(0,\tfrac{1}{2})$ is arbitrary, letting first $\Lambda\to\infty$ and then $\epsilon\downarrow 0$ yields
$\lim_{\Lambda\to+\infty}\Bigl(\Lambda^{-\frac{n}{2s}}(\ln\Lambda)^{\frac{n}{2s}}{\mathcal{N}}(\Lambda)\Bigr)=\frac{n+2s}{2s}A_{s}=(2\pi)^{-n}s^{\frac{n}{2s}}\omega_{n}\,|\Omega|.$
Taking $\Lambda=\lambda_{k}^{s+\operatorname*{Log}}(\Omega)$, implies that
(4.7) $\lim_{k\to\infty}k\,\big(\lambda_{k}^{s+\operatorname*{Log}}(\Omega)\big)^{-\frac{n}{2s}}\big(\ln\lambda_{k}^{s+\operatorname*{Log}}(\Omega)\big)^{\frac{n}{2s}}=(2\pi)^{-n}s^{\frac{n}{2s}}\omega_{n}\,|\Omega|=:C_{0}.$
Directly from (4.7), for any $\delta\in(0,1)$ there exists $k_{\delta}$ such that for all $k\geq k_{\delta}$,
$(1-\delta)C_{0}\leq k\,\big(\lambda_{k}^{s+\operatorname*{Log}}(\Omega)\big)^{-\frac{n}{2s}}\big(\ln\lambda_{k}^{s+\operatorname*{Log}}(\Omega)\big)^{\frac{n}{2s}}\leq(1+\delta)C_{0}.$
Equivalently,
$\Big(\frac{k}{(1+\delta)C_{0}}\Big)^{\frac{2s}{n}}\leq\frac{\lambda_{k}^{s+\operatorname*{Log}}(\Omega)}{\ln\lambda_{k}^{s+\operatorname*{Log}}(\Omega)}\leq\Big(\frac{k}{(1-\delta)C_{0}}\Big)^{\frac{2s}{n}}.$
In particular, $\lambda_{k}^{s+\operatorname*{Log}}(\Omega)\to\infty$, and the above bounds imply $\lambda_{k}^{s+\operatorname*{Log}}(\Omega)\sim k^{\frac{2s}{n}}\ln k$, hence
$\frac{\ln\lambda_{k}^{s+\operatorname*{Log}}(\Omega)}{\ln k}\longrightarrow 1\qquad\text{as }k\to\infty.$
Therefore, dividing the previous inequality by $\ln k$ and using $\ln\lambda_{k}^{s+\operatorname*{Log}}(\Omega)\sim\ln k$, we obtain
$\lim_{k\to\infty}\frac{\lambda_{k}^{s+\operatorname*{Log}}(\Omega)}{k^{\frac{2s}{n}}\ln k}=C_{0}^{-\frac{2s}{n}}.$
Finally, since $C_{0}=(2\pi)^{-n}s^{\frac{n}{2s}}\omega_{n}|\Omega|$, we compute
$C_{0}^{-\frac{2s}{n}}=\Big((2\pi)^{-n}s^{\frac{n}{2s}}\omega_{n}|\Omega|\Big)^{-\frac{2s}{n}}=(2\pi)^{2s}\,s^{-1}\,(\omega_{n}|\Omega|)^{-\frac{2s}{n}},$
and hence
$\lim_{k\to\infty}\frac{\lambda_{k}^{s+\operatorname*{Log}}(\Omega)\,k^{-\frac{2s}{n}}}{\ln k}=\frac{2s}{n}\,C_{0}^{-\frac{2s}{n}}=\frac{2}{n}\,(2\pi)^{2s}\,(\omega_{n}|\Omega|)^{-\frac{2s}{n}}.$
The proof ends. ∎